### v1

# Cross-lingual Offensive Language Detection: A Systematic Review of Datasets, Transfer Approaches and Challenges

Abstract.
The growing prevalence and rapid evolution of offensive language in social media amplify the complexities of detection, particularly highlighting the challenges in identifying such content across diverse languages. This survey presents a systematic and comprehensive exploration of Cross-Lingual Transfer Learning (CLTL) techniques in offensive language detection in social media. Our study stands as the first holistic overview to focus exclusively on the cross-lingual scenario in this domain. We analyse 67 relevant papers and categorise these studies across various dimensions, including the characteristics of multilingual datasets used, the cross-lingual resources employed, and the specific CLTL strategies implemented. According to “what to transfer”, we also summarise three main CLTL transfer approaches: instance, feature, and parameter transfer. Additionally, we shed light on the current challenges and future research opportunities in this field. Furthermore, we have made our survey resources available online, including two comprehensive tables that provide accessible references to the multilingual datasets and CLTL methods used in the reviewed literature.
Disclaimer: Due to the nature of this work, some examples may contain offensive text and hate speech.
1. Introduction
The prevalence of offensive language on social media platforms, such as Facebook and Twitter, has become a pressing concern in recent years. This surge can be attributed to the anonymity provided by these platforms and the lack of strict regulations and effective mechanisms to restrain such behaviour (Fortuna and Nunes, 2018). While social media platforms have fostered connections and bridged global distances, they have in turn enabled the spread of hate speech and various forms of offensive language (Chhabra and Vishwakarma, 2023). Offensive language broadly refers to expressions that may upset or annoy others, while hate speech escalates beyond mere offensive behaviour to inciting discrimination, hostility, or violence against individuals or groups based on their race, religion, ethnicity, gender, or other identity factors (for Peace, [n. d.]). Hate speech is considered a human rights violation in many legal systems due to its potential to cause real-world harm and perpetuate discrimination against targeted groups (UNSPAHS, 2019). Hence, there is a growing scholarly focus on analysing and developing approaches to mitigate hate speech in addition to offensive content.
The Natural Language Processing (NLP) community has proposed a variety of research techniques to detect offensive content based on both traditional machine learning and advanced neural network-based approaches. While early work on offensive language detection focused on monolingual settings (i.e. English), research in recent years has increasingly focused on tackling the problem in multilingual settings to broaden the ability to deal with more languages, a key requirement given the presence of diverse languages in social media platforms (Pamungkas et al., 2021b). This however generally hindered by the limited availability of labelled data and the high variability of offensive language across diverse cultures and languages (Plaza-del arco et al., 2023).
Cross-Lingual Transfer Learning (CLTL) emerges as a promising direction to mitigate challenges associated with data scarcity by leveraging domain knowledge from high- to low-resource languages (Pikuliak et al., 2021). Given its adaptability and compatibility with neural networks, CLTL has found successful applications across NLP tasks (Pan and Yang, 2010). Some pioneering efforts have also integrated CLTL for offensive language detection in low-resource languages (Jiang and Zubiaga, 2021; Shi et al., 2022; Zhou et al., 2023). However, open challenges remain in building effective and generalisable cross-lingual models and understanding linguistic gaps.
The earliest work in our survey is from 2018 (Mathur et al., 2018b) and, since then, interest in cross-lingual offensive language detection has increased, but is yet to be covered in a comprehensive survey. Existing surveys concentrate on monolingual detection or specific characteristics of offensive language, while the unique CLTL application in offensive language detection is seldom considered. Relevant surveys exist (Pamungkas et al., 2021b; Chhabra and Vishwakarma, 2023), but their scope goes beyond cross-lingual scenarios, also covering cross-domain and multimodal dimensions, and hence lacking our survey’s focus on CLTL.
1.1. Scope of the Survey
Our objective is to systematically review existing literature, elucidating different CLTL techniques employed in the task of offensive language detection. Consequently, this survey discusses 67 papers in this area. This is the first systematic and holistic overview of recent studies, challenges and open research avenues of using CLTL to detect offensive language in social media across languages. We describe those studies according to multiple aspects, including the multilingual datasets employed, cross-lingual resources leveraged, levels of transfer, and cross-lingual strategies applied. Inspired by Pikuliak et al. (2021), we outline three CLTL transfer approaches based on the level at which knowledge transfer occurs across languages, namely instance transfer, feature transfer, and parameter transfer. We also illustrate prevalent transfer strategies utilised based on these transfer levels. In addition, we highlight challenges and posit potential opportunities for future research in this field. Furthermore, we make our survey resources available online,111https://github.com/aggiejiang/crosslingual-offensive-language-survey including two comprehensive tables that provide accessible references to the multilingual datasets and CLTL methods used in surveyed papers to facilitate easy comparison and discovery of related works.
1.2. Survey Structure
The rest of the paper is organised as follows. Section 2 describes the definitions of offensive language and the framework of cross-lingual detection, and distinguishes ‘cross-lingual’ from related terminologies such as ‘multilingual’. Related surveys are also summarised. Section 3 summarises our survey methodology. Then we summarise and analyse the multilingual datasets leveraged in surveyed papers in Section 4. In Section 5, we describe diverse linguistic resources and tools used in cross-lingual studies. Section 6 discusses the three transfer levels of CLTL in offensive language detection, complemented by transfer strategies based on these layers. Section 7 presents current challenges and future research directions in this area, concluding the survey in Section 8.
2. Background and Related Work
2.1. Cross-linguality in offensive Language Detection
2.1.1. Task Definition and Workflow
The task of cross-lingual hate speech detection focuses on the identification and categorisation of offensive or hateful text across different languages. The primary objective for CLTL is to leverage the knowledge acquired from one language (source language with more resources) to enhance the hate speech detection in another language (target language with less resources), especially when the labelled data in the target language is scarce or unavailable (Pamungkas et al., 2021b). The workflow of cross-lingual hate speech is presented below:
Data Preparation. Let and represent the source and target languages. Then let and represent the source and target datasets respectively, where denotes text data and denotes labels indicating that the text is offensive or not. The cross-lingual task contains a predictive function , a feature space , and a label space , where 0 denotes non-offensive and 1 denotes offensive content.
Cross-lingual Training. The training stage for cross-lingual transfer can use only or both and . For only as training data, a model is trained to learn a function , where is the feature space of . Then adapt the learned function to a function , where is the feature space of , with minimal loss in performance. For both and as training data, a model is trained to learn a function , where is the joint feature space of . Various cross-lingual training strategies could be employed for cross-lingual transfer to obtain , where more details are described in Section 6.
Model Adaptation. If label is available for , fine-tune on full or few shots of to obtain a model , further adapting the function to the target domain: .
Detection on Target Language. Apply predictive functions or to to detect offensive content in and predict labels: or .
2.1.2. Cross-lingual & Multilingual & Code-mixing
While “cross-lingual learning” typically denotes the application of CLTL techniques, the term “multilingual learning” is sometimes used interchangeably (Pikuliak et al., 2021). Some multilingual data may contain code-mixed content, i.e. a mixture of languages in the same sentence (Ibrohim and Budi, 2019a). Although some researchers draw parallels between these terms and the relationship between transfer learning and multi-task learning (Wang and Zheng, 2015), we adopt a broader perspective on cross-lingual learning. We define cross-lingual learning as the process of transferring knowledge between different languages, a concept that inherently subsumes multilingual learning.
2.2. Related Surveys
We found 18 surveys covering offensive language detection from different angles.
General Offensive Language Detection. Surveys on offensive language detection provide general insights into the field (Schmidt and Wiegand, 2017; Fortuna and Nunes, 2018; Jahan and Oussalah, 2023; Mishra et al., 2019). Key works by Schmidt and Wiegand (2017) and Fortuna and Nunes (2018) stand out as the two most cited reviews in the NLP community, providing definitions of hate speech and related terms, summarising previous methods, and discussing the existing limitations and challenges. Some of these surveys focus on in-depth discussions of specific aspects when reviewing studies in hate speech detection, such as data sources (Alkomah and Ma, 2022; Vidgen and Derczynski, 2020; Poletto et al., 2021), methodologies (Alkomah and Ma, 2022; Ayo et al., 2020; Demilie and Salau, 2022), and challenges (Ayo et al., 2020; Kiritchenko et al., 2021; Yin and Zubiaga, 2021; Garg et al., 2023). Additionally, several surveys have also delved into related areas like cyberbullying (Yi and Zubiaga, 2023) and the intersection of hate speech and fake news (Demilie and Salau, 2022), and others review ethical considerations in the development of hate speech detection (Vidgen and Derczynski, 2020; Kiritchenko et al., 2021).
Cross-lingual and Multilingual Offensive Language Detection. Several surveys focus on studies in single but non-English languages, such as Arabic (Husain and Uzuner, 2021; Al-Hassan and Al-Dossari, 2019), Indian (Dowlagar and Mamidi, 2021) and Ethiopian (Demilie and Salau, 2022) languages. Several surveys for general hate speech detection also provide a brief summary of multilingual or cross-lingual contributions in this area (Poletto et al., 2021; Al-Hassan and Al-Dossari, 2019; Yin and Zubiaga, 2021). However, systematic reviews focusing explicitly on cross-lingual and multilingual perspectives remain scarce (Pamungkas et al., 2021b; Chhabra and Vishwakarma, 2023). Pamungkas et al. (2021b) provide an overview of available datasets and approaches in cross-domain and cross-lingual settings. Chhabra and Vishwakarma (2023) conduct a comprehensive review of multimodal and multilingual hate speech detection, including benchmark datasets, state-of-the-art methods, challenges, and the latest evaluation techniques.
Our survey instead focuses on cross-lingual and multilingual offensive language detection by providing a systematic review into available datasets, cross-lingual resources, existing CLTL approaches, related challenges and future research directions.
3. Survey Methodology
We adopt the structured PRISMA framework (Page et al., 2021) for our systematic review to ensure both replicability and transparency. A detailed flowchart of our search protocol is provided in Figure 1, which illustrates the progression of publications through different phases of our selection process.
3.1. Selection Process
Search Databases
We retrieve publications from scientific databases including ACL Anthology,222https://aclanthology.org/ Google Scholar,333https://scholar.google.com/ and DBLP Computer Science Bibliography.444https://dblp.org/
Keyword Selection
We create a list of 99 pairwise combinations of words in (i) and (ii): (i) 9 keywords related to offensive language, such as “hate”, “abus*”, “offens*”, “toxic”, “cyberbully”, “gender”, “sexis*”, “misogyn*”, “rac*”; (ii) 11 keywords related to multilingual and cross-lingual scenarios, such as “cross-lingual”, “multilingual”, “low resource”, “under resource”, “cross language”, “between language”, “cross linguistic”, “transfer”, “code switch”, “zero shot”, “few shot”.
Search Process
In addition to studies matching the 99 keyword pairs above, we also identify and include 5 additional studies cited in relevant surveys. This search was conducted on July 29th, 2023.
Eligibility Criteria
Following removal of duplicates and irrelevant papers, the remainder 455 publications were checked against the inclusion / exclusion criteria in Table 1.
Include | Exclude |
(i) Describe cross-lingual approaches to detect general abusive content (toxic, hate, etc.) |
(i) Study transfer across dimensions other than languages (e.g. platforms, domains, datasets) |
(ii) Describe cross-lingual approaches to detect subtypes of abusive content (e.g. sexism, racism) |
(ii) Study low-resource languages without knowledge transfer approaches |
(iii) Describe multilingual approach with knowledge transfer between languages |
(iii) Relevant studies from competition reports and papers |
(iv) Study text classification cases in cross-lingual hate speech scenarios |
(iv) Master or PhD dissertations |
3.2. Summary of Surveyed Studies
In the end we obtained 67 relevant papers, with their temporal distribution shown in Figure 2 and showing a growing trend from the first publication in 2018 to 2022.
4. Multilingual Hate Speech Datasets
We investigate 82 datasets utilised in existing surveyed studies on cross-lingual hate speech detection across different aspects. We summarise these datasets and their characteristics in Table 2.
Ref | Year | Topic | Source | Language | Size | Label | #Cit | CM? | Avail? |
(Paul et al., 2023) |
2023 |
cyberbullying |
en, hi |
22k |
1 |
¡50 |
(Jeong et al., 2022) |
2022 |
offense |
NAVER-news, Youtube |
ko |
40,429 |
1,2,3 |
¡10 |
(Jiang et al., 2022) |
2022 |
Sexism |
zh |
8,969 |
1,2,3 |
¡50 |
(Salaam et al., 2022) |
2022 |
offense |
en, fr, de, es |
37,221 |
1 |
¡10 |
(Deng et al., 2022) |
2022 |
offense |
Weibo, Zhihu |
zh |
37,480 |
1,2 |
¡50 |
(Gupta et al., 2022) |
2022 |
abuse |
ShareChat |
hi, kn, ml, ta, te |
92,881 |
1 |
¡10 |
(Chakravarthi et al., 2022) |
2022 |
offense |
Youtube |
ta |
60k |
1,3 |
¡50 |
(Arango Monnar et al., 2022) |
2022 |
offense |
es |
9,834 |
2 |
¡10 |
(Bhattacharya et al., 2020) |
2022 |
aggressiveness, misogyny |
Youtube |
bn |
25k |
1 |
¡100 |
(Kumar and Sachdeva, 2022) |
2022 |
cyberbullying |
Facebook, Twitter |
en, hi |
6,500 |
1 |
¡50 |
(Pavlopoulos et al., 2021) |
2021 |
toxicity |
Civil Comments |
en |
10,629 |
1,2 |
¡100 |
(Burtenshaw and Kestemont, 2021) |
2021 |
toxicity |
Ask.fm |
nl |
10,189 |
2 |
¡10 |
(Hande et al., 2021b) |
2021 |
offense |
Youtube |
kn, ml, ta |
71,691 |
1,3 |
¡20 |
(Mathew et al., 2021) |
2021 |
hate speech |
Gab, Twitter |
en |
20,148 |
1,2 |
¡500 |
(Romim et al., 2021) |
2021 |
hate speech |
Facebook, Youtube |
bn |
30k |
1 |
¡50 |
(Modha et al., 2021) |
2021 |
hate speech, offense |
en, hi, mr |
13,755 |
1,4 |
¡100 |
(Amjad et al., 2022) |
2021 |
abuse |
ur |
8,400 |
1 |
¡50 |
(Vidgen et al., 2020) |
2021 |
hate speech |
open-source platform |
en |
41,255 |
1,2 |
¡100 |
(Gaikwad et al., 2021) |
2021 |
offense |
mr |
2499 |
1,2,3 |
¡50 |
(Khan et al., 2021) |
2021 |
hate speech, offense |
roman ur |
5k |
1 |
¡50 |
(Mubarak et al., 2021) |
2021 |
offense |
ar |
10k |
1,2 |
¡250 |
(Pitenis et al., 2020) |
2020 |
offense |
el |
4,779 |
1 |
¡250 |
(Charitidis et al., 2020) |
2020 |
hate speech |
en, fr, de, el, es |
264,035 |
1 |
¡50 |
(Glavaš et al., 2020) |
2020 |
hate speech, abuse |
Facebook, Fox news, Twitter, Wikipedia |
sq, hr, en, de, ru, tr |
109,955 |
1 |
¡50 |
(Shekhar et al., 2020) |
2020 |
offense |
24sata, Eesti Ekspress, Veˇcernji List |
hr, et |
62.6m |
¡10 |
(Akhter et al., 2020) |
2020 |
offense, abuse |
Youtube |
Roman ur, ur |
12,171 |
1 |
¡100 |
(Rizwan et al., 2020) |
2020 |
hate speech, offense |
Roman ur |
10,012 |
1,2 |
¡50 |
(Mandl et al., 2021) |
2020 |
offense |
Twitter, Youtube |
en, de, hi, ml, ta |
15,047 |
1,2 |
¡250 |
(Sanguinetti et al., 2020) |
2020 |
hate speech |
it |
12,081 |
1 |
¡100 |
(Kennedy et al., 2020) |
2020 |
hate speech |
Gab |
en |
27,655 |
1 |
¡250 |
(Çöltekin, 2020) |
2020 |
offense |
tr |
36,232 |
1,3 |
¡250 |
(Kumar et al., 2020) |
2020 |
aggressiveness |
Youtube |
bn, en, hi |
15k |
1 |
¡250 |
(Sigurbergsson and Derczynski, 2020) |
2020 |
abuse, offense |
Facebook, Reddit, Twitter |
da |
3,600 |
1,3 |
¡250 |
(Hande et al., 2020) |
2020 |
offense |
Youtube |
kn |
7,671 |
1,3 |
¡100 |
(Zampieri et al., 2020) |
2020 |
offense |
ar, da, en, el, tr |
9m |
1,3 |
¡500 |
(Ljubešić et al., 2019) |
2019 |
offense |
en, sl |
22,877 |
1,2 |
¡50 |
(Arata, 2019) |
2019 |
cyberbullying |
ja |
4,096 |
1 |
¡10 |
(Arango et al., 2019) |
2019 |
hate speech |
en |
14,949 |
1 |
¡250 |
(Ibrohim and Budi, 2019b) |
2019 |
abuse, hate speech |
id |
5,561 |
1,2,3 |
¡250 |
(Mulki et al., 2019) |
2019 |
abuse, hate speech, toxicity |
at |
5,846 |
1 |
¡250 |
(Pereira-Kohatsu et al., 2019) |
2019 |
hate speech |
es |
6k |
1 |
¡100 |
(Fortuna et al., 2019) |
2019 |
hate speech, offense |
pt |
5,668 |
1 |
¡100 |
(Poletto et al., 2021) |
2019 |
cyberbullying, toxicity |
pl |
11,041 |
1,2 |
¡50 |
(Strus̈ et al., 2019) |
2019 |
offense |
de |
7,025 |
1,2 |
¡250 |
(Basile et al., 2019) |
2019 |
aggressiveness |
en |
19,600 |
1,3 |
¡750 |
(Zampieri et al., 2019) |
2019 |
offense |
en, es |
14k |
1,3 |
¡700 |
(Mandl et al., 2019) |
2019 |
hate speech, offense |
Twitter, Facebook |
en, de, hi |
17,657 |
1,2,3 |
¡100 |
(Chung et al., 2019) |
2019 |
hate speech |
en, fr, it |
4,078 |
1,2 |
¡250 |
(Ousidhoum et al., 2019) |
2019 |
hate speech |
ar, en, fr |
13,014 |
1,2,3 |
¡250 |
(Chiril et al., 2019) |
2019 |
sexism, misogyny |
fr |
3,085 |
1 |
¡50 |
(Bosco et al., 2018) |
2018 |
hate speech |
Facebook, Twitter |
en, it |
4k |
1 |
¡250 |
(de Gibert et al., 2018) |
2018 |
hate speech |
Stormfront |
en |
10,568 |
1 |
¡500 |
(Ptaszyński et al., 2018) |
2018 |
cyberbullying |
Formspring.me |
en |
12,772 |
1 |
¡50 |
(Alakrot et al., 2018) |
2018 |
abuse, offense |
Youtube |
ar |
167,549 |
1 |
¡100 |
(Bohra et al., 2018) |
2018 |
hate speech |
en, hi |
4,575 |
1 |
¡250 |
(ElSherief et al., 2018) |
2018 |
hate speech |
en |
27,330 |
1,2 |
¡250 |
(Fersini et al., 2018a) |
2018 |
misogyny |
en, it |
20k |
1,2,3 |
¡250 |
(Mathur et al., 2018a) |
2018 |
abuse, hate speech, offense |
en, hi |
17,698 |
1,2 |
¡100 |
(Kumar et al., 2018) |
2018 |
aggressiveness |
Facebook, Twitter |
en, hi |
39k |
1,2,3 |
¡250 |
(Mathur et al., 2018b) |
2018 |
offense |
en, hi |
3,679 |
1,2 |
¡250 |
(Sanguinetti et al., 2018) |
2018 |
aggressiveness, hate speech, offense |
it |
6k |
2,4 |
¡250 |
(Founta et al., 2018) |
2018 |
abuse, hate speech |
en |
80k |
1,2 |
¡550 |
(Wiegand et al., 2018) |
2018 |
hate speech, offense |
de |
8,541 |
1,2 |
¡500 |
(Fersini et al., 2018b) |
2018 |
misogyny |
en, es |
8,115 |
1,2,3 |
¡250 |
(Golbeck et al., 2017) |
2017 |
harassment, offense, racism |
BlockTogether, Twitter |
en |
35k |
1 |
¡100 |
(Devyatkin et al., 2017) |
2017 |
offense |
ru |
493 |
1,2 |
¡50 |
(Gao and Huang, 2017) |
2017 |
hate speech |
Fox news |
en |
1,528 |
1 |
¡250 |
(Wulczyn et al., 2017) |
2017 |
hate speech |
Wikipedia |
en |
115,737 |
1 |
¡750 |
(de Pelle and Moreira, 2017) |
2017 |
hate speech, offense |
g1.globo.com |
pt |
1,250 |
1 |
¡100 |
(Ross et al., 2017) |
2017 |
hate speech, offense |
de |
541 |
1,2,4 |
¡500 |
(Alfina et al., 2017) |
2017 |
hate speech |
id |
713 |
1 |
¡250 |
(Fortuna, 2017) |
2017 |
hate speech |
pt |
5,668 |
1,2 |
¡50 |
(Bretschneider and Peters, 2017) |
2017 |
hate speech, offense |
de |
5,836 |
1,3 |
¡100 |
(Davidson et al., 2017) |
2017 |
offense |
en |
24,802 |
1,2 |
¡2200 |
(Mubarak et al., 2017) |
2017 |
abuse, hate speech, offense |
ar |
32k |
1 |
¡500 |
(Waseem and Hovy, 2016) |
2016 |
hate speech |
en |
16,914 |
1 |
¡1550 |
(Van Hee et al., 2015) |
2015 |
cyberbullying |
Ask.fm |
nl |
85,462 |
2,4 |
¡250 |
(Dybala et al., 2010) |
2010 |
cyberbullying |
informal websites of Japanese secondary schools |
ja |
2,999 |
1 |
¡100 |
Topics
The topic distribution among datasets shows that “offensiveness” and “hate speech” are the two most frequently addressed topics, accounting for 34.23% and 32.43% respectively. Some cover general offense, such as abusiveness, toxicity, cyberbullying, harassment and aggressiveness. In addition, some datasets narrow their scope to targeted issues, like sexism, misogyny and racism.
Data Sources
Most studies tend to collect data from openly accessible social media sources such as Twitter and Facebook. Our review of data sources reveals a pronounced reliance on popular social media platforms, with Twitter notably encompassing 47.19% of the datasets. This is followed by Youtube and Facebook, which amount to 11.24% and 10.11% respectively. While these mainstream platforms dominate, news websites like Fox News and NAVER news, and open forums like Reddit and 2ch, also contribute to the multilingual datasets. And there is a rich diversity in the data source origins. Sources like Weibo (Chinese), 2ch (Russian), g1.globo (Brazilian Portuguese) and Eesti Ekspress (Estonian) represent some non-English speaking regions as more culturally specific platforms. Some datasets originate from platforms known for controversial content or specific user bases, such as Stormfront and alt-right websites.
Languages and Families
Datasets cover 32 languages across 10 language families, where one dataset may contain more than one language or family. We show the distribution of languages and their language families in Figure 3. The distribution of languages in the datasets highlights a representative language of English, accounting for 25.95% of the data, followed by Hindi at 9.16%, and German at 6.87%. Notably, the datasets exhibit a relatively balanced representation across several languages, such as Spanish, Arabic, Italian, French, Portuguese and Turkish, each ranging from around 2% to 5%. This distribution reveals a strong emphasis on Indo-European languages in hate speech studies, complemented by a notable presence of Afro-Asiatic languages (represented primarily by Arabic), with other language families being less represented and restricting cross-lingual research. Some datasets (14.6%) are also dedicated to code-mixed content.
Dataset Size
Over half of the datasets, accounting for 51.20%, contain - instances. This is closely followed by 35.40% of datasets that have instances between and . On the other hand, large datasets are rare, with only one dataset’s size exceeding , and two between and . We can observe that the majority of datasets fall within the smaller size ranges, which indicates the challenges in collecting and manually annotating large-scale labelled datasets, as well as model generalisability for cross-lingual hate speech research.
Data Labelling and Distributions
According to diverse annotation schemes, four different types of labels are often employed: (i) binary labels, (ii) fine-grained categories of offensive content, (iii) attack targets, and (iv) intensity scores. A majority of datasets, 76 in total, utilise the simpler binary labelling, of which 36 solely provide binary labels. There are also datasets combining binary labels with either fine-grained categories (20), attack targets (9), or intensity scores (1). A smaller number of datasets (10) combine binary labels, fine-grained categories, and attack targets. Interestingly, only four datasets consider intensity scores, suggesting that quantifying the severity of offensive content is less common in current research.
Dataset Availability
We find 66 out of 82 datasets (80%) are publicly accessible, with the other 16 unavailable, in some cases with the possibility of contacting the authors to request the data.
Competitions and Shared Tasks
The rising concern surrounding online abuse has led to the establishment of numerous competitions and shared tasks within both national (e.g., GermEval, Evalita, IberLEF) and international (e.g., SemEval) evaluation campaigns. These open scientific competitions release benchmark datasets and invite participants to submit detection results and detailed system reports (Poletto et al., 2021). Table 3 lists 24 such competitions and shared tasks from 2018 to 2023, each addressing different facets of offensive content across multiple languages.
Prominent topics include offense and hate speech. Specifically, GermEval competitions in 2018 (Wiegand et al., 2018) and 2019 (Strus̈ et al., 2019) centre around identifying offensive content in German tweets. OffensEval, initially in 2019 focused only on English content (Zampieri et al., 2019), expanded to include multiple languages in the subsequent edition in 2020 (Zampieri et al., 2020). The task in OSACT4 (Mubarak et al., 2020), the 4th Workshop on Open-Source Arabic Corpora and Processing Tools, delves into the challenges of offensive language detection in Arabic, especially its dialectal forms. HaSpeeDe has consistently addressed hate speech, with its iterations focusing on various general and particular hate speech topics (Bosco et al., 2018; Sanguinetti et al., 2020; Lai et al., 2023). For instance, HaSpeeDe (Bosco et al., 2018) and HaSpeeDe 2 (Sanguinetti et al., 2020) concentrate on hate speech against immigrants and Muslims, while HaSpeeDe 3 (Lai et al., 2023) explores hate speech in strongly polarised debates, in particular concerning political and religious topics. Similarly, the HASOC series, recurring annually from 2019 to 2023, delves into hate speech and offensive content identification across various languages (Mandl et al., 2019, 2021; Modha et al., 2021; Satapara et al., 2023). TRAC workshops, held in 2018 (Kumar et al., 2018) and 2020 (Kumar et al., 2020), tackled trolling, aggression, and cyberbullying. Other competitions, like PolEval (Ptaszynski et al., 2019) and UrduThreat (Amjad et al., 2022), focus on specific languages (such as Polish and Urdu) or topics (e.g., cyberbullying and threatening languages). The Jigsaw Multilingual Toxic Comment Classification competition on Kaggle attracted over 1600 participants to build multilingual models using English-only training data. Another Kaggle competition, IIIT-D Multilingual Abusive Comment Identification, focuses on abusive comments in Indic languages. Additionally, some tasks focus on more specific topics, such as sexism and misogyny. AMI is organised by IberEval (Fersini et al., 2018b) and Evalita (Fersini et al., 2018a) in 2018, specifically addressing Spanish and Italian misogynous content respectively. EXIST is held for three consecutive years from 2021 to 2023 and aims to identify sexism in social networks (Rodríguez-Sánchez et al., 2021, 2022; Plaza et al., 2023).
The majority of these benchmark datasets originates from social media platforms like Twitter, and often conduct annotations that go beyond simple binary labels to finer-grained ones, enabling to further and broaden research in cross-lingual hate speech detection.
Name | Description | Year | Topic | Language | #Teams |
Identification of Offensive Language |
2018 |
offense |
de |
20 | |
AMI@IberEval (Fersini et al., 2018b) |
Automatic Misogyny Identification |
2018 |
misogyny |
en, es |
11 |
AMI@Evalita (Fersini et al., 2018a) |
Automatic Misogyny Identification |
2018 |
misogyny |
en, it |
16 |
HaSpeeDe@Evalita (Bosco et al., 2018) |
Hate Speech Detection |
2018 |
hate speech |
it |
9 |
Aggression Identification of Hindi-English Code-mixed Data |
2018 |
aggressiveness |
en, hi |
30 | |
HatEval@SemEval (Basile et al., 2019) |
Multilingual Detection of Hate Speech Against Immigrants and Women in Twitter |
2019 |
hate speech |
en, es |
74 |
HASOC@FIRE (Mandl et al., 2019) |
Offensive Content Identification in Indo-European Languages |
2019 |
hate speech, offense |
en, hi, de |
37 |
Task2@GermEval (Strus̈ et al., 2019) |
Identification of Offensive Language |
2019 |
offense |
de |
13 |
Task6@PolEval (Ptaszynski et al., 2019) |
Automatic Cyberbullying Detection in Polish Twitter |
2019 |
cyberbullying |
pl |
9 |
OffensEval@SemEval (Zampieri et al., 2020) |
Multilingual Offensive Language Identification in Social Media |
2020 |
offense |
en, ar, da, el, tr |
145 |
HASOC@FIRE (Mandl et al., 2021) |
Offensive Language Identification in Tamil, Malayalam, Hindi, English and German |
2020 |
hate speech, offense |
en, hi, de, ta, ml |
40+ |
HaSpeeDe@Evalita (Sanguinetti et al., 2020) |
Hate Speech Detection |
2020 |
hate speech |
it |
14 |
Aggression and Gendered Aggression Identification |
2020 |
aggressiveness |
en, hi, bn |
19 | |
Jigsaw Multilingual Toxic Comment Classification |
2020 |
toxicity |
en, es, tr, pt |
1621 | |
Arabic Offensive Language Detection |
2020 |
offense |
ar |
27 | |
HASOC@FIRE (Modha et al., 2021) |
Offensive Content Identification in English and Indo-Aryan Languages and Conversational Hate Speech |
2021 |
hate speech, offense |
en, hi, mr |
65 |
EXIST@IberLEF (Rodríguez-Sánchez et al., 2021) |
Sexism Identification in Social Networks |
2021 |
sexism |
en, es |
31 |
UrduThreat@FIRE (Amjad et al., 2022) |
Abusive and Threatening Language Detection in Urdu |
2021 |
abuse, threat |
ur |
19 |
Moj Multilingual Abusive Comment Identification across Indic Languages |
2021 |
abuse |
10+ Indic |
54 | |
HASOC@FIRE (Satapara et al., 2023) |
Offensive Content Identification in English and Indo-Aryan Languages |
2022 |
hate speech, offense |
en, hi, de, mr |
12 |
EXIST@IberLEF (Rodríguez-Sánchez et al., 2022) |
Sexism Identification in Social Networks |
2022 |
sexism |
en, es |
19 |
HaSpeeDe@Evalita (Lai et al., 2023) |
Political and Religious Hate Speech Detection |
2023 |
hate speech |
it |
6 |
EXIST@IberLEF (Plaza et al., 2023) |
Sexism Identification in Social Networks |
2023 |
sexism |
en, es |
28 |
Hate Speech and Offensive Content Identifica- |
2023 |
hate speech, |
en, hi, de, |
tion in English and Indo-Aryan Languages |
offense |
Indo-Aryan |
5. Cross-lingual Resources
5.1. Multilingual linguistic Resources
There are some basic but essential linguistic resources that facilitate the task of cross-lingual hate speech detection across two or more languages. Among these, multilingual lexicons and parallel corpora stand out as foundational resources, frequently utilised by researchers to bridge linguistic gaps and enhance model performance in CLTL.
Multilingual Lexicons (word-aligned): contain words, phrases, or terms in two or more languages. These lexicons often provide direct translations or equivalents of terms across the languages they cover, such as HurtLex.555https://github.com/valeriobasile/hurtlex
Parallel Corpora (sentence-aligned): refer to datasets that consist of texts in two or more languages, where each text in one language has a direct translation in the other languages.
Due to the lack of sufficient labelled datasets for hate speech, multilingual lexicons can help bridge this gap by providing connections between resource-rich and resource-poor languages (Pamungkas et al., 2021a; Jiang and Zubiaga, 2021), while parallel corpora provide translations of labelled instances from resource-rich to resource-poor languages, to facilitate CLTL (Rivaling, 2021; Stamou et al., 2022; Bigoulaeva et al., 2023). Cross-lingual word embeddings are also created by utilising these multilingual resources. In addition, these are also valuable linguistic resources in other NLP tasks, such as machine translation and multilingual model training.
Multilingual linguistic resources are most often created for specific domains, such as hate speech and abusive language. While hate speech patterns can be universal, they can also be specific to certain cultures or languages. These resources provide insights into how hate speech manifests differently across languages and cultures, furthering the understanding of cultural nuances (Nozza, 2021).
5.2. Machine Translation and Transliteration Tools
Translation refers to converting content from one language into another, whereas transliteration is the process of converting text from one script into another, in both cases preserving the meaning of the original text (Hande et al., 2021b). Machine translation tools can translate datasets from one language to another, such as Google Translate,666https://translate.google.com Microsoft Translator,777https://translator.microsoft.com/ DeepL,888https://www.deepl.com/en/translator) and machine translation models (e.g., mBART (Liu et al., 2020) and mT5 (Xue et al., 2021)). Besides, machine transliteration tools are essential for code-mixed datasets, such as Google Transliteration,999https://www.google.co.in/inputtools/try/ Microsoft Transliteration,101010https://www.microsoft.com/en-us/translator/business/translator-api/ AI4Bharat transliteration application,111111https://github.com/AI4Bharat/IndianNLP-Transliteration and transliteration python packages (e.g., indic-transliteration121212https://github.com/indic-transliteration/indic_transliteration_py).
However, in offensive language detection, some translation tools may lead to translation errors or cultural nuances that are inconsistent with the original meaning (Pamungkas et al., 2021a). For instance, derogatory terms or slurs without direct equivalents in other languages, or their severity might differ across cultures. Additionally, idiomatic expressions that convey hate or abuse in one language might lose their offensive meaning when translated literally. Therefore, translation quality is crucial to maintain the effectiveness and robustness of cross-lingual detection models.
5.3. Multilingual Representations
Multilingual representations, as language independent representations, are often used in cross-lingual transfer learning. With these representations we can directly address the difference between source and target languages by projecting the text into a shared feature space. Then it is easier to project the behaviour of the model. In general, we can distinguish between using word-level representations and sentence-level representations.
Multilingual distributional representations encode words from multiple languages in a single distributional word vector space. Cross-lingual transfer of word embeddings aims to establish the semantic mappings among words in different languages by learning the transformation functions over the corresponding word embedding spaces. In this space, semantically similar words (e.g. ‘cat’ in English and ‘katze’ in German) are close together independently of the language. Some prominent distributional embeddings are Multilingual GloVe (Ferreira et al., 2016), Multilingual FastText (Grave et al., 2018), Babylon (Smith et al., 2017), and Multilingual Unsupervised or Supervised word Embeddings (MUSE) (Lample et al., 2018b, a). During the training, multilingual distributional representations usually require additional cross-lingual resources, e.g., bilingual dictionaries or parallel corpora.
Multilingual contextualised representations on sentence level work on a similar principle, but they use sentences instead of words for alignments. They are usually based on an auto-encoder architecture, in which the model is pushed to create similar and context-aware representations for parallel sentences, such as Language Agnostic Sentence Representations (LASER) (Artetxe and Schwenk, 2019) and Language Agnostic BERT Sentence Embeddings (LabSE) (Feng et al., 2022).
5.4. Monolingual and Multilingual Pre-trained Language Models
Pre-trained language models (PLMs) are a SOTA technique in NLP, as a language model trained on large amounts of data. PLMs can be initialised and further trained or fine-tuned with target data for cross-lingual transfer learning. The best-known examples of PLMs are Embeddings from Language Models (ELMo) (Peters et al., 2018) and Bidirectional Encoder Representations from Transformers (BERT) (Devlin et al., 2019).
Multilingual PLMs are an extension of monolingual PLMs, trained on multiple languages at the same time. They learn to understand the connections between languages by developing a shared linguistic representation. There is no need to provide multilingual PLMs with any additional information about interlanguage relations, and cross-lingual transfer is possible between any languages (Pikuliak et al., 2021). Prominent examples include Multilingual BERT (mBERT) (Devlin et al., 2019) and XLM-RoBERTa (XLM-R) (Conneau et al., 2020), which have demonstrated remarkable capabilities in understanding and processing multiple languages simultaneously.
5.5. Language-Agnostic Resources
Language-agnostic resources (e.g., emojis, capitals, and punctuations) can be seen as common traits among different languages, sharing similar knowledge to enhance the semantic connection across languages. For instance, emojis are generally associated with emotional expressions, which in turn are associated with various forms of online hate (Corazza et al., 2020).
6. Cross-lingual Transfer Approaches
In this section, we systematically describe and compare different cross-lingual techniques to offensive language online. Referring to the way of identifying different transfer learning techniques in Pan and Yang (2010) and Pikuliak et al. (2021), we analyse cross-lingual approaches in all reviewed papers and ultimately decide to categorise them according to “what to transfer” – the level at which knowledge transfer or sharing occurs. The overall hierarchy of various cross-lingual transfer approaches is depicted in Figure 4. Cross-lingual transfer paradigms consist of three main categories:
Instance Transfer: Instances are transferred on the data level between source and target languages, including texts or labels transfer.
Feature Transfer: Linguistic knowledge is shared or transferred on the feature level between source and target languages.
Parameter Transfer: Parameter values are transferred between language models. This effectively transfers the behaviour of the model from source to target languages.
We look at transfer levels in Table 4 and models in Table 5. Additionally, a comprehensive Table is provided online131313https://github.com/aggiejiang/crosslingual-offensive-language-survey to summarise all techniques utilised in these studies.
Transfer Level | Reference |
Instance |
(Mathur et al., 2018b; Ibrohim and Budi, 2019a; Chiril et al., 2019; Pamungkas and Patti, 2019; Pant and Dadu, 2020; Arango et al., 2022; Elouali et al., 2020; Pamungkas et al., 2020; Aluru et al., 2020; Jiang and Zubiaga, 2021; Song et al., 2021; Hande et al., 2021b; Vashistha and Zubiaga, 2021; Pamungkas et al., 2021a; Aluru et al., 2021; Salaam et al., 2022; Ryzhova et al., 2022; Boulouard et al., 2022; Shi et al., 2022; Deshpande et al., 2022; Zia et al., 2022; Das et al., 2022; zahra El-Alami et al., 2022; Kumar and Sachdeva, 2022; Zhou et al., 2023) |
Feature |
(Sohn and Lee, 2019; Corazza et al., 2020; Bansal et al., 2020; Aluru et al., 2020; Arango et al., 2021; Hahn et al., 2021; Rodríguez et al., 2021; Vitiugin et al., 2021; Pelicon et al., 2021a; Bigoulaeva et al., 2021, 2022; De la Peña Sarracén and Rosso, 2022; Shi et al., 2022; Jiang and Zubiaga, 2023; Bigoulaeva et al., 2023; Paul et al., 2023) |
Parameter |
(Mathur et al., 2018b; Chiril et al., 2019; Kapoor et al., 2019; Sohn and Lee, 2019; Pamungkas and Patti, 2019; Ousidhoum et al., 2019; Pant and Dadu, 2020; Arango et al., 2022; Elouali et al., 2020; Stappen et al., 2020; Glavaš et al., 2020; Pamungkas et al., 2020; Ranasinghe and Zampieri, 2020; Aluru et al., 2020; Markov et al., 2021; Burtenshaw and Kestemont, 2021; Firmino et al., 2021; Tiţa and Zubiaga, 2021; Jiang and Zubiaga, 2021; Rodríguez et al., 2021; Song et al., 2021; Pelicon et al., 2021b, a; Bigoulaeva et al., 2021; Ranasinghe and Zampieri, 2021a; Hande et al., 2021a, b; Ranasinghe and Zampieri, 2021b; Vashistha and Zubiaga, 2021; Ranasinghe and Zampieri, 2021c; Nozza, 2021; Gaikwad et al., 2021; Pamungkas et al., 2021a; Aluru et al., 2021; Salaam et al., 2022; Husain and Uzuner, 2022; Gupta et al., 2022; Montariol et al., 2022; Ryzhova et al., 2022; Muti and Barrón-Cedeño, 2022; Phung and Cloos, 2022; Bigoulaeva et al., 2022; Boulouard et al., 2022; R”̈ottger et al., 2022a; De la Peña Sarracén and Rosso, 2022; Shi et al., 2022; Deshpande et al., 2022; Eronen et al., 2022; Mozafari et al., 2022; Zia et al., 2022; R”̈ottger et al., 2022b; Das et al., 2022; zahra El-Alami et al., 2022; Kumar and Sachdeva, 2022; Plaza-del arco et al., 2023; Zhou et al., 2023; Bigoulaeva et al., 2023; Awal et al., 2023) |
Type of Model | Model | Reference |
Machine Learning |
NB |
(Ibrohim and Budi, 2019a) |
Based |
LR |
(Aluru et al., 2020; Vashistha and Zubiaga, 2021; Aluru et al., 2021; Arango Monnar et al., 2022; Deshpande et al., 2022; Paul et al., 2023) |
SVM |
(Ibrohim and Budi, 2019a; Pamungkas and Patti, 2019; Pamungkas et al., 2020; Jiang and Zubiaga, 2021; Rodríguez et al., 2021; Bigoulaeva et al., 2021; Paul et al., 2023) |
DT | ||
RF |
(Ibrohim and Budi, 2019a; Burtenshaw and Kestemont, 2021; Rodríguez et al., 2021) |
other |
(Hahn et al., 2021) |
Deep Learning |
MLP |
(Jiang and Zubiaga, 2021; Pelicon et al., 2021a; Kumar and Sachdeva, 2022; Paul et al., 2023) |
Based |
CNN |
(Mathur et al., 2018b; Elouali et al., 2020; Markov et al., 2021; Jiang and Zubiaga, 2021; Bigoulaeva et al., 2021, 2022; Jiang and Zubiaga, 2023; Bigoulaeva et al., 2023; Paul et al., 2023) |
LSTM |
(Chiril et al., 2019; Kapoor et al., 2019; Pamungkas and Patti, 2019; Ousidhoum et al., 2019; Arango et al., 2022; Stappen et al., 2020; Pamungkas et al., 2020; Markov et al., 2021; Burtenshaw and Kestemont, 2021; Jiang and Zubiaga, 2021; Vitiugin et al., 2021; Bigoulaeva et al., 2021; Pamungkas et al., 2021a; Phung and Cloos, 2022; Bigoulaeva et al., 2022; Shi et al., 2022; Arango Monnar et al., 2022; Kumar and Sachdeva, 2022; Bigoulaeva et al., 2023; Paul et al., 2023) |
GNN |
(De la Peña Sarracén and Rosso, 2022) |
other |
(Ousidhoum et al., 2019; Bansal et al., 2020; Jiang and Zubiaga, 2021; Vashistha and Zubiaga, 2021; Arango Monnar et al., 2022; Deshpande et al., 2022; Kumar and Sachdeva, 2022) |
Transformer Based |
BERT |
(Sohn and Lee, 2019; Pamungkas et al., 2020; Aluru et al., 2020; Markov et al., 2021; Arango et al., 2021; Hande et al., 2021a; Vashistha and Zubiaga, 2021; Muti and Barrón-Cedeño, 2022; Boulouard et al., 2022; Arango Monnar et al., 2022; Zia et al., 2022; zahra El-Alami et al., 2022; Plaza-del arco et al., 2023; Zhou et al., 2023; Jiang and Zubiaga, 2023; Paul et al., 2023) |
RoBERTa |
(Pant and Dadu, 2020; Hande et al., 2021a; Zia et al., 2022; Plaza-del arco et al., 2023; Zhou et al., 2023) |
DistilBERT |
(Hande et al., 2021a) |
AraBERT |
(Husain and Uzuner, 2022; Boulouard et al., 2022; zahra El-Alami et al., 2022) |
ULMFiT | ||
mBERT |
(Sohn and Lee, 2019; Glavaš et al., 2020; Aluru et al., 2020; Burtenshaw and Kestemont, 2021; Tiţa and Zubiaga, 2021; Jiang and Zubiaga, 2021; Rodríguez et al., 2021; Song et al., 2021; Pelicon et al., 2021b, a; Hande et al., 2021b; Ranasinghe and Zampieri, 2021c; Nozza, 2021; Pamungkas et al., 2021a; Aluru et al., 2021; Montariol et al., 2022; Muti and Barrón-Cedeño, 2022; Bigoulaeva et al., 2022; Boulouard et al., 2022; De la Peña Sarracén and Rosso, 2022; Deshpande et al., 2022; Eronen et al., 2022; Das et al., 2022; zahra El-Alami et al., 2022; Zhou et al., 2023; Bigoulaeva et al., 2023; Awal et al., 2023) |
XLM |
(Corazza et al., 2020; Stappen et al., 2020; Hande et al., 2021a; Zhou et al., 2023) |
XLM-R |
(Pant and Dadu, 2020; Glavaš et al., 2020; Ranasinghe and Zampieri, 2020; Firmino et al., 2021; Tiţa and Zubiaga, 2021; Jiang and Zubiaga, 2021; Song et al., 2021; Ranasinghe and Zampieri, 2021a; Hande et al., 2021a, b; Ranasinghe and Zampieri, 2021b, c; Nozza, 2021; Gaikwad et al., 2021; Salaam et al., 2022; Gupta et al., 2022; Montariol et al., 2022; Ryzhova et al., 2022; De la Peña Sarracén and Rosso, 2022; Arango Monnar et al., 2022; Eronen et al., 2022; Mozafari et al., 2022; Zia et al., 2022; Plaza-del arco et al., 2023; Awal et al., 2023) |
XLM-T | ||
MuRIL | ||
other |
(Rodríguez et al., 2021; Pelicon et al., 2021b; Hande et al., 2021a, b; Husain and Uzuner, 2022; Gupta et al., 2022; Montariol et al., 2022; Plaza-del arco et al., 2023; Jiang and Zubiaga, 2023) |
Large Language |
FLAN-T5 |
(Plaza-del arco et al., 2023) |
Model |
mT0 |
(Plaza-del arco et al., 2023) |
6.1. Instance Transfer
Instances in the cross-lingual hate speech task are comprised of two elements – texts and corresponding labels – in both source and target languages. Although the single source language data is not directly applicable in cross-lingual transfer, there are certain parts of the instance that can still be repurposed together with target language data via instance transfer. Instance transfer (also called instance projection) revolves around the transfer of specific data elements (either text or label information) between source and target languages, which is a key technique in the realm of cross-lingual offensive language detection.
To implement instance transfer, a correspondence between instances in the source and target languages must be created on the data level. Correspondence refers to the pair of instances with the same meaning of texts or identical labels (Pikuliak et al., 2021), e.g., parallel data possess the same label matching, and translated text retains the same label of its original counterpart. Based on the established correspondence, texts or labels can be transitioned from one language to the other . Subsequently, these projected texts or labels can be used for further training and fine-tuning stages.
Pairs of instances in source and target languages that share identical labels are referred to as corresponding instances. All data-based transfer techniques necessitate the presence of corresponding texts/labels or a mechanism to generate them. Techniques for instance transfer are introduced distinctly for both text and label levels. For label-based transfer, the primary strategies include annotation projection and pseudo-labelling. Meanwhile, For text-based transfer, the predominant approaches are machine translation and text alignment.
6.1.1. Label-Based Transfer
Annotation Projection
It involves leveraging parallel corpora to transfer annotations from a source language to a target language. Parallel corpora consist of sentence-aligned source and target language data, ensuring higher quality than machine-generated translations. Since data in the source language is paired with target language data, labels in the source language can be directly projected onto the target language. For example, if a sentence in the source language is labelled as hateful, its corresponding text in the target language can be inferred to have the same label. It can effectively create labelled data for the target language without manual annotation, especially useful when building datasets for languages where annotating from scratch might be challenging due to linguistic complexities or lack of expert annotators.
Pseudo Labelling
It is a semi-supervised learning technique, where a model trained on labelled source data is used to make predictions on unlabeled target data (Lee et al., 2013). High-confidence predictions are treated as “pseudo-labels” for the unlabelled target data. By using high-confidence predictions from a model trained in the source language, it is able to generate labelled target data in another language. This augmented target data can then be used to train or fine-tune models for offensive language detection in the target language. To generate pseudo labels, some studies use an ensemble-based approach to generate labels based on majority voting for unlabelled target datasets, and use bootstrapped target dadasets to further fine-tune the trained model with source training samples (Bigoulaeva et al., 2021, 2022, 2023). Hande et al. (2021b) directly use a pre-trained multilingual language model (i.e. XLM-R (Conneau et al., 2020)) to predict the pseudo labels. Zia et al. (2022) firstly fine-tune a pre-trained XLM-R on gold-labeled source language data, and then use it to create a new psuedo-labelled dataset in the target language.
6.1.2. Text-Based Transfer
Machine Translation
In the absence of sufficient annotated parallel data resources, machine translation provides an efficient alternative strategy to achieve text-level transmission. In general, it leverages translation tools (Pamungkas et al., 2020, 2021a) or models (Glavaš et al., 2020) to translate labelled datasets between source and target languages, thereby augmenting the training data. Labels for translated data are the same as those in the original labelled dataset. Translation can be either directed or undirected: (i) target to source (Ibrohim and Budi, 2019a; Arango et al., 2022; Boulouard et al., 2022; Aluru et al., 2020, 2021); (ii) source to target (Glavaš et al., 2020; Song et al., 2021; Salaam et al., 2022); (iii) translate both source and target to each other (Sohn and Lee, 2019; Pamungkas and Patti, 2019; Jiang and Zubiaga, 2021; zahra El-Alami et al., 2022; Das et al., 2022). In addition, back-translation is commonly used as a data augmentation technique in cross-lingual scenarios by using various transformations on the original source data to create new samples for further model training (Ouyang et al., 2021). Back-translation refers to translating a source sentence to a target language and then reverting it back to the original language, aiming to generate paraphrased instances of the original texts.
Text Alignment
Similar to machine translation, existing mapping techniques between languages are employed to generate labelled data for the low-resource language. Shi et al. (2022) begin by using a shared space between two language vectors to identify the most similar sentence in the target language to the labelled data in the source language. The identified sentences are then treated as labelled data for the target language, and assigned the same labels of the source data. Ryzhova et al. (2022) augment training datasets by generating attacked source datasets. They apply an adversarial attack algorithm (Jin et al., 2020) to the source sentence by only replacing words that the model considers important, ensuring that the new sentence is semantically similar to the old one.
6.2. Feature Transfer
Feature transfer approaches transform linguistic features to aid in the detection of offensive language. Rather than translating source texts or projecting source labels, they focus on extracting salient linguistic features from source and target languages, and aligning them into a shared feature space, retaining the essence of the text consistent across languages.
Cross-lingual Word Embeddings (CLWEs), also known as Multilingual Word Embeddings (MWEs), are commonly used for feature transfer in offensive language detection. They train monolingual word embeddings (such as Word2Vec or FastText) on multiple languages, yielding vectors that capture semantic similarities relying on shared representations between languages and thus enabling representations that are common across languages.
In cross-lingual offensive language detection, various approaches utilise existing CLWEs or retrofit CLWEs. We discuss three main techniques: multilingual distributional representations, multilingual contextualised representations, and retrofitting word embeddings.
6.2.1. Multilingual Distributional Representations
Pre-trained distributional word embeddings are most commonly used as embedding layers in neural networks for different language data inputs. Among these, MUSE embeddings have emerged as a predominant choice. They are extensively utilised to extract multilingual features and are often integrated with both traditional machine learning models, such as Gradient Boosted Decision Trees (Arango et al., 2022), and advanced deep learning architectures (i.e. LSTM (Pamungkas and Patti, 2019; Pamungkas et al., 2020, 2021a; Arango et al., 2021; Burtenshaw and Kestemont, 2021), BiLSTM (Bigoulaeva et al., 2021, 2022, 2023; Jiang and Zubiaga, 2021), CNN-GRU (Deshpande et al., 2022; Aluru et al., 2021, 2020), Capsule Networks (Jiang and Zubiaga, 2021), Sluice networks (Ousidhoum et al., 2019)), and BERT (Pamungkas et al., 2020)). Other notable multilingual distributional vectors, such as Multilingual GloVe and FastText, have also been employed in architectures like BiLSTM (Chiril et al., 2019) and Capsule Networks (Jiang and Zubiaga, 2021). In addition, Babylon multilingual embeddings have been paired with Sluice networks for multilingual hate speech detection (Ousidhoum et al., 2019). Bansal et al. (2020) integrate bilingual switching features into the Hierarchical Attention Network (HAN) architecture, enhancing the transfer knowledge for cross-lingual detection.
Furthermore, some studies have ventured into cross-lingual projection using the MUSE mapping method, aligning monolingual FastText embeddings in English and German to produce their own CLWEs (Bigoulaeva et al., 2021, 2022, 2023). Paul et al. (2023) align English and Hindi FastText embeddings by performing Canonical Correlation Analysis (CCA),141414https://www.mathworks.com/help/stats/canoncorr.html and project them into a shared vector space where they are maximally correlated. Kapoor et al. (2019) use multilingual datasets to train embeddings in order to capture specific distributional representations of tweets. De la Peña Sarracén and Rosso (2022) propose a graph auto-encoder framework to learn embeddings of a set of texts in an unsupervised way, adding language-specific knowledge via Universal Sentence Encoder (USE).
6.2.2. Multilingual Contextualised Representations
Multilingual contextualised representations, typically derived at the sentence level from PLMs (such as mBERT (Devlin et al., 2019) and XLM-R (Conneau et al., 2020)), can capture deeper semantic and contextual relationships between words and phrases. Some researchers have leveraged hidden features with richer semantic information from the first embedding layer of mBERT (Arango et al., 2021; Pamungkas et al., 2021a; Rodríguez et al., 2021), DistilmBERT (Vitiugin et al., 2021; Hande et al., 2021b), XLM (Stappen et al., 2020), XLM-R (Hande et al., 2021b) as feature extraction layers. These embeddings are either extracted as standalone embedding layers (Arango et al., 2021; Pamungkas et al., 2021a; Vitiugin et al., 2021; Rodríguez et al., 2021) or utilised as frozen embedding layers (Stappen et al., 2020; Hande et al., 2021b) in deep neural networks, such as BiLSTM, yielding enhanced performance over traditional distributional embeddings. Multilingual sentence embeddings LASER have also been used across various models, including Support Vector Machine (Rodríguez et al., 2021), Logistic Regression (Aluru et al., 2021, 2020; Pamungkas et al., 2021a; Deshpande et al., 2022), Random Forest (Arango et al., 2021; Rodríguez et al., 2021), XGBoost (Arango et al., 2021), Multi-Layer Perceptron (Pelicon et al., 2021a) and LSTM (Vitiugin et al., 2021). Rodríguez et al. (2021) propose to use LabSE and mBERT representations with SVM-based and tree-based classifiers, achieving improved performance.
6.2.3. Retrofitting Word Embeddings
Retrofitting pre-trained word embeddings by infusing multilingual domain knowledge stands as a promising strategy to amplify the semantic relationships across languages, thereby enhancing the efficacy and scalability of models in cross-lingual offensive language detection. Pant and Dadu (2020) employ a supervised FastText model trained on the sarcasm detection corpus (Swami et al., 2018) to improve the identification of hate speech content in English and Hindi languages. Arango et al. (2021) construct hate speech-specific word embeddings by aligning monolingual Word2Vec embeddings using hate-specific bilingual dictionaries (such as HurtLex), which are able to capture non-traditional translations of words between languages. Additionally, Hahn et al. (2021) learn semantic subspace-based representations to model profane languages on both word and sentence levels. Jiang and Zubiaga (2023) propose a cross-lingual domain-aware semantic specialisation system to construct sexism-aware word embeddings. They retrofit pre-trained FastText word vectors by integrating in-domain and out-of-domain linguistic knowledge (such as lexico-semantic relations) into the specialised feature space.
6.2.4. Infusing Additional Features
Additional external features are commonly used in identifying hate speech across languages, such as domain-specific (Pamungkas and Patti, 2019; Hahn et al., 2021), language-specific (Zhou et al., 2023; Arango et al., 2021), and typographic features (Kumar and Sachdeva, 2022; Markov et al., 2021). Specifically, domain-specific features are essential to assist the hate-related knowledge transfer. The multilingual lexicon, HurtLex, has been employed in multiple studies (Pamungkas et al., 2020, 2021a; Pamungkas and Patti, 2019; Jiang and Zubiaga, 2021). Some researchers also utilise hateful word pairs to construct domain-aware or semantic-based representations across languages (Jiang and Zubiaga, 2023; Hahn et al., 2021; Arango et al., 2021), due to the lack of hate-specific patterns in general-purpose multilingual embeddings. In addition, infusing language-specific features can help integrate linguistic and cultural knowledge for geographically sensitive tasks. It can be in different forms, such as language switching pattern matrix (Bansal et al., 2020), bilingual pairs (Jiang and Zubiaga, 2023), cross-cultural similarities (Arango et al., 2021), and social dynamics among users (Nozza, 2021). This feature can provide a richer understanding of cultural dimensions in hateful content (Zhou et al., 2023). Furthermore, typographic features provide a language-agnostic and domain-agnostic perspective to help identify online hate due to their shared meanings across languages, which includes capitals, punctuations, and emojis (Kumar and Sachdeva, 2022; Markov et al., 2021; Corazza et al., 2020). Stylometric features also remain persistent with respect to language variations, as the writing style of toxic content can be correlated to emotional profile of social media users (Markov et al., 2021).
6.3. Parameter Transfer
Performance of NLP models is dependent on the values of model parameters. Parameter transfer assumes that some parameters or prior distributions of hyperparameters are transferred between different languages within one model or between individual models. Most parameter transfer approaches rely on inductive transfer learning in cross-lingual offensive language detection (Pan and Yang, 2010). Inspired by Pikuliak et al. (2021), we outline three distinct scenarios for all surveyed studies, depending on how the source and target data are utilised during the process of parameter transfer, namely zero-shot transfer, joint learning, and cascade learning (see Figure 5). Only source data is used for training in zero-shot transfer, while both source and target data are applied during the training process in joint learning and cascade learning. Joint learning utilises both simultaneously to train the model, but cascade learning leverages them in the different stages of training.
Parameter-level transfer strategies can be sophisticated and diverse in terms of their model architectures and training procedures, incorporating one or more of the transfer scenarios above. Table 6 has summarised the possible correlations between these transfer strategies and their application scenarios, and will discuss more details in the following subsections regarding scenarios and strategies utilised on parameter-level transfer.
Transfer Strategy | Zero-Shot Transfer | Joint Learning | Cascade Learning |
Multilingual PLMs |
Domain Adaptation |
Multi-source Training |
Multi-channel Learning |
Multi-task Learning |
Meta Learning |
Ensemble Learning |
6.3.1. Transfer Scenarios
We discuss three different scenarios of parameter transfer.
Zero-Shot Transfer
It is a promising scenario for cross-lingual offensive language detection when labelled data in the target language is scarce or lacking, by enabling detection in a language not seen during training. The model is trained on labelled source data from one or more languages not including the target language, and is then tested on the unseen target language. This is sometimes called direct or model transfer. We found 25 papers under this transfer paradigm, including subtypes:
Single Source to Single Target: Considered a type of domain adaptation (see Section 6.3.2), models are trained on labelled data from a single source language, then used on an unseen target language (Glavaš et al., 2020; Firmino et al., 2021; Tiţa and Zubiaga, 2021; Rodríguez et al., 2021; Pelicon et al., 2021a; Nozza, 2021; Montariol et al., 2022; Eronen et al., 2022; Zhou et al., 2023; Gupta et al., 2022; Ranasinghe and Zampieri, 2021a, b; Gaikwad et al., 2021; Muti and Barrón-Cedeño, 2022; Das et al., 2022; Plaza-del arco et al., 2023). It is the most frequently observed (62.5%) of zero-shot transfer strategies in our survey.
Multiple Sources to Single Target: Models are trained on labelled data from multiple source languages (excluding the target language) and then evaluated on a single target language (Ryzhova et al., 2022; Montariol et al., 2022; Aluru et al., 2020). It is a type of multi-source training strategy where the combination of training data excludes the target (see Section 6.3.2).
Pseudo-Target Augmented Training: The source data is translated into the parallel target data via machine translation, and then models are trained on both the original source data and the translated target data together (Stappen et al., 2020) or in parallel (Pamungkas and Patti, 2019; Pamungkas et al., 2021a, 2020; Jiang and Zubiaga, 2021). It falls under the multi-channel learning strategy when using both source and translated target data in parallel151515Pseudo-Target Augmented Training has overlaps with the joint learning scenario. It distinguishes itself by using pseudo target texts for training instead of real target texts. and fusing them before the output layer of the model. This allows the model to learn nuances specific to the target language from the pseudo content.
Parameter Frozen: A model is trained on source data first, and then a subset of its parameters (e.g., weights from specific layers) is saved to initialise another model for target language evaluation (Ranasinghe and Zampieri, 2020, 2021c; Gaikwad et al., 2021; Phung and Cloos, 2022). It is a type of domain adaptation,161616Parameter frozen in zero-shot transfer scenario shares similarities with the cascade learning scenario. However, while cascade learning continues to fine-tune the trained model with target data, zero-shot transfer directly applies the trained model to test the target data without further fine-tuning. and captures language knowledge from certain layers of the source model to provide a head start for the target model, potentially leading to faster convergence and better generalisation.
It is worth highlighting that these methods rely on multilingual PLMs, especially often used for Single Source to Single Target and Multiple Sources to Single Target scenarios. mBERT and XLM-R are two of the most prominent multilingual PLMs in surveyed papers, because of their popularity, widespread adoption and SOTA performance across multilingual NLP tasks. They together account for 80.8% of papers in the zero-shot transfer scenario. Additionally, some papers release domain- and language-specific PLMs in a multilingual setting (Gupta et al., 2022), choose other multilingual PLMs like XLM (Stappen et al., 2020) and language-specific MuRIL (Gupta et al., 2022; Das et al., 2022), or combine multilingual word embeddings with monolingual neural models like LSTM (Pamungkas and Patti, 2019; Jiang and Zubiaga, 2021) and BERT (Pamungkas et al., 2020).
Joint Learning
It typically involves training a model between languages based on multiple channels, tasks or objectives simultaneously, or jointly assimilating multiple types of data or modalities. The model might have some parameters that are shared across languages as well as some that are specific to each, so it is also called parameter sharing (Pikuliak et al., 2021). The transfer of knowledge happens only on shared parameters. Changes to the shared parameters can affect each other between languages. They are interrelated and can provide complementary information.
In this scenario, both source and target data are used at the same time during the training stage. Parameters can be shared between source and target channels within one model or between individual source and target models. We found 23 papers in this group. The distinction among these strategies mainly lies in the stage of knowledge transfer across languages and the extent of parameter sharing (some or all parameters shared) (Kumar and Sachdeva, 2022). Based on these criteria, we categorise them into three primary fusion stages (see Figure 6):
Early fusion: A combination of datasets or embeddings from both source and target languages, with the entire model sharing parameters. It is commonly adopted in few-shot learning (Stappen et al., 2020; Gaikwad et al., 2021; Firmino et al., 2021; Das et al., 2022; Gupta et al., 2022; Zhou et al., 2023) and multi-source training (Muti and Barrón-Cedeño, 2022; Salaam et al., 2022; Ranasinghe and Zampieri, 2021a; Deshpande et al., 2022; Vashistha and Zubiaga, 2021; Elouali et al., 2020; Das et al., 2022; Gupta et al., 2022; Markov et al., 2021; Chiril et al., 2019), where datasets from the target language are incorporated during training. In few-shot learning, the model is trained by utilising a very small amount of target samples combined with a more extensive source dataset, while multi-source training typically takes a combined dataset of source and target languages or integrates multiple datasets including multiple non-target datasets together with the target dataset for training.
Model-level fusion: A concatenation of high-level hidden features within a model, and only specific layers or components of the model share parameters, facilitating a more nuanced transfer of knowledge. It commonly occurs in parallel model architectures, such as multi-channel learning (Sohn and Lee, 2019; Kumar and Sachdeva, 2022), multi-task learning (Hande et al., 2021a; Ousidhoum et al., 2019) and meta-learning (Mozafari et al., 2022; Awal et al., 2023).
Late fusion: A combination of different channels or models before the output layer, and only the parameters of the output layer for predictions are shared, ensuring a final consolidation of knowledge from various languages, often used in multi-channel architectures (Ibrohim and Budi, 2019a).
Early fusion is the most prevalent strategy due to its straightforward integration of source and target data for better generalisation, while late fusion provides a modular approach, allowing for the independent training of individual models that can later be merged. However, unlike early fusion, model-level fusion effectively overcomes the curse of high dimensionality and the synchronisation demands among diverse features. Moreover, it ensures that interactions between different languages are not isolated as in late fusion (Kumar and Sachdeva, 2022), showing its robustness and resilience.
Cascade Learning
In the cascade learning scenario, both source and target data are used at different times during the training. First, an existing PLM is directly used or a new model is trained with source data, and then the trained model is fine-tuned with target data, where subsets of parameters are shared between different stages of training and fine-tuning. Different strategies for training and fine-tuning have been used in 19 of our surveyed papers.
Many use existing SOTA pre-trained multilingual models serving as the foundation,171717Cascade learning fine-tunes PLMs using target data, while zero-shot transfer fine-tunes PLMs using non-target data. while some studies train a model with one or multiple source datasets. Subsequently, the fine-tuning stage in cascade learning is applied across an entire (Kapoor et al., 2019; Firmino et al., 2021; Aluru et al., 2021; Pelicon et al., 2021b; Husain and Uzuner, 2022) or only a limited subset of target data (i.e. few-shot learning) (Ranasinghe and Zampieri, 2021a; Pelicon et al., 2021b), optimising the model adaptability. R”̈ottger et al. (2022a) indicate that an initial training phase on source data (English) could increase model performance when there is little fine-tuning data in the target language, where source data can partly substitute target data in few-shot settings. Some researchers delve into consistent fine-tuning (Ryzhova et al., 2022; Mozafari et al., 2022; Awal et al., 2023) or iterative fine-tuning (i.e. meta learning) (Mozafari et al., 2022; Awal et al., 2023), performing multiple rounds of fine-tuning on the PLMs to further enhance its ability to transfer knowledge.
Additionally, given the limited availability of target data, data augmentation techniques have emerged. They aim to enrich the target dataset to enable fine-tuning, by enhancing label diversity or expanding the text corpus. Some studies have attempted to generate pseudo labels for unlabelled target datasets (Bigoulaeva et al., 2021, 2022, 2023; Zia et al., 2022) (see Section 6.1.1), while others augment and refine pseudo texts by exploring advanced techniques, such as word alignment to project samples (Shi et al., 2022), adversarial algorithm to generate attack samples (Ryzhova et al., 2022), and domain-specific target data filtering (Glavaš et al., 2020).
6.3.2. Hybrid Transfer Strategies
According to the utilisation of source and target data, these strategies can be broadly covered in one or more parameter transfer scenarios mentioned in Section 6.3.1. That is, these strategies are not mutually exclusive. A single scenario can seamlessly integrate multiple transfer strategies (see Table 6).
Multilingual Pre-Trained Language Models
These models are regarded as a cornerstone of parameter-level transfer techniques for low-resource scenarios, especially in cross-lingual offensive language detection. They are trained on large-scale multilingual corpora, learning commonalities and differences between languages. Therefore, a single PLM can share the same underlying sub-word vocabulary and semantic representations across languages (Devlin et al., 2019). Multilingual PLMs can be initialised as the base model with pre-trained parameters, then fine-tuned on available hate speech data in all three scenarios of parameter transfer (zero-shot transfer, joint learning and cascade learning) (Muti and Barrón-Cedeño, 2022; Firmino et al., 2021). Beyond this, they can also employed in the other two transfer levels (data and feature). They are able to augment target datasets by predicting pseudo labels for unlabelled data and generating synthetic target texts. And they can be used as multilingual representations, where subsequent models are constructed based on these source representations (Rodríguez et al., 2021).
Domain Adaptation
It involves taking a model trained (or pre-trained) in the source language and adapting its parameters to initialise a new model for the low-resource language, aiming to leverage source data to improve performance on target data (Lee et al., 2021; Glavaš et al., 2020). Its advantage is that it can capture offensive language knowledge and linguistic similarities between the languages within the model’s parameters. In the zero-shot transfer scenario, parameters of the trained model are entirely or partially used on the target data for prediction (Gaikwad et al., 2021; Ranasinghe and Zampieri, 2021c). In addition, cascade learning adopts a more subtle approach. That is, the entire or specific subsets of the model’s parameters (like specific layers) are frozen to retain their original state (Ranasinghe and Zampieri, 2021a), while others are fine-tuned to better align with the target data, often using a smaller target dataset (as few-shot learning) (Pelicon et al., 2021b; R”̈ottger et al., 2022a).
Multi-Source Training
This approach often entails the amalgamation of labelled data from multiple source languages and potentially the target language in diverse ways. By synthesising information across multiple languages, this strategy is able to capture the diversities and commonalities of linguistic patterns, demonstrating significant efficacy in the cross-lingual setting. It is extensively utilised across all three transfer scenarios, serving as an augmented cross-lingual learning approach during different stages of training or fine-tuning. In multi-source training, the most prevalent way is to train or fine-tune the model on multiple source languages. This can be executed in the training phase of joint learning (R”̈ottger et al., 2022b; Muti and Barrón-Cedeño, 2022; Salaam et al., 2022; Deshpande et al., 2022; Vashistha and Zubiaga, 2021; Elouali et al., 2020; zahra El-Alami et al., 2022; Das et al., 2022; Hande et al., 2021a; Markov et al., 2021; Chiril et al., 2019) or the fine-tuning phase of cascade learning (Ranasinghe and Zampieri, 2021a; Gupta et al., 2022) in conjunction with the target language, or in isolation from the target language for zero-shot transfer (Ryzhova et al., 2022; Montariol et al., 2022). Deshpande et al. (2022) offer a more focused perspective by narrowing the linguistic scope to specific language families. By training models on languages within one language family, they then assess the model’s performance on each member of that family, revealing the interplay of linguistic knowledge within related languages. It is worth noting that merging training examples from different languages can be detrimental to cross-lingual performance in cases where offensive language domains are too distant (Glavaš et al., 2020).
Multi-Channel Learning
This strategy typically refers to the use of multiple types of input representations or sources concurrently for a single task, where multiple embeddings or pre-processed versions of the text are often channelled as parallel input streams. In this architecture, each channel processes the input independently through parallel layers. These independent processing channels are eventually combined based on their outputs for subsequent learning or classification stages. Such a design ensures that diverse input representations are holistically integrated, capturing a broader spectrum of linguistic knowledge across languages. Multi-channel learning finds its applications mainly in the zero-shot transfer and joint learning scenarios. When applied to cross-lingual offensive language detection, this approach often treats datasets from different languages as distinct input channels. Each channel adopts different neural networks or PLMs based on diverse input languages, and these channels are then merged into the final classification stage (Kumar and Sachdeva, 2022; Stappen et al., 2020; Sohn and Lee, 2019). A significant improvement to this strategy is the incorporation of machine translation. By translating labelled source data to pseudo target data, low-resource target data is enriched and a bridge is constructed between source and target languages, enhancing knowledge relationships and a more seamless transfer of languages (Pamungkas and Patti, 2019; Pamungkas et al., 2020, 2021a; Jiang and Zubiaga, 2021; Ibrohim and Budi, 2019a).
Multi-Task Learning
A joint learning environment training a model on several interrelated tasks simultaneously, where the proficiency learned from one task can boost performance on another. A key point of this strategy is the shared representation of hidden features across specific layers of the model. Such shared layers enable different tasks to mutually benefit from common representations, while also preserving their distinctiveness through task-specific output layers. The training process is holistic, aiming to optimise the model’s performance across all tasks. This is typically achieved by incorporating the loss functions from each task. Multi-task learning emerges as a powerful method for cross-lingual offensive language detection within the parameter-level transfer techniques across three transfer scenarios. Models can be trained on offensive language detection and a wide range of related auxiliary tasks, such as Sentiment Analysis, Named Entity Recognition (NER), Dependency Parsing, and Part-Of-Speech (POS) Tagging (Montariol et al., 2022; Hande et al., 2021a; Ousidhoum et al., 2019). Additionally, they can be fine-tuned on data in various offensive categories, such as aggressive content (Ranasinghe and Zampieri, 2020, 2021c). This multi-task training strategy enables the model to identify and process offensive content with higher accuracy and enhances the model’s ability to grasp linguistic nuances across languages.
Meta Learning
This strategy, often referred to as “learning to learn”, has become a popular few-shot learning technique in NLP. It involves training models to learn the optimal initialisation of parameters, allowing them to be fine-tuned with a small amount of data from the target language. This technique finds its application within the joint learning and cascade learning transfer scenarios. It enables rapid adaptation to new unseen languages, making it especially useful in scenarios where labelled data is limited. The main meta learning methods include optimisation-based (Finn et al., 2017) and metric-based techniques (Koch et al., 2015). Two papers in our review employ optimisation-based meta learning framework. Montariol et al. (2022) study meta learning for the problem of few-shot hate speech detection in low-resource languages. They propose a cross-lingual meta learning-based approach based on optimisation-based Model-Agnostic Meta-Learning (MAML) and Proto-MAML models, fine-tuning the base learner XLM-R with parallel few-shot datasets in different target languages. Awal et al. (2023) propose HateMAML, a model-agnostic meta-learning-based framework that uses a semi-supervised self-refinement strategy to fine-tune a better PLM for unseen data in target language, showing effective performance for hate speech detection in low-resource languages.
Ensemble Learning
It leverages multiple models or “learners” to collaboratively make decisions. Instead of relying on the output of a single model, ensemble learning combines all predictions from multiple models to produce a final prediction based on majority voting (Paul et al., 2023; Markov et al., 2021). It takes advantage of the different strengths of each model, mitigating the weaknesses and biases of any single model and avoiding overfitting, improving robustness and reliability over a single classifier. This is particularly beneficial in cross-lingual settings, where language biases can pose challenges to individual models. In cross-lingual offensive language detection, ensemble learning is mainly applied in the zero-shot, joint learning, and cascade learning transfer scenarios. Deep ensemble models have emerged as the predominant choice in recent studies (Bigoulaeva et al., 2021, 2022, 2023; Paul et al., 2023), while the integration of PLMs within ensemble learning frameworks is also common (Burtenshaw and Kestemont, 2021; Ryzhova et al., 2022; Markov et al., 2021). Each model in the ensemble can be trained on different languages or linguistic features, allowing the ensemble model to capture diverse linguistic patterns and reduce test errors for target languages (Ryzhova et al., 2022).
6.4. Summary of Cross-lingual Approaches
For transfer scenarios, zero-shot transfer does not use any target data, while joint and cascade learning do. With zero-shot learning, we always transfer the whole model, while with joint and cascade learning we might transfer only a subset of parameters. Some overlaps of transfer strategies might exist between each transfer scenario because multiple strategies can be conducted based on more than one transfer scenario.
Model distributions. We provide an overview of diverse models and their frequency used in recent cross-lingual studies, which reveals a clear trend towards more advanced architectures. Traditional machine learning-based models have been employed 24 times, while deep learning-based models have seen a slightly higher usage (38 times) due to their complex structure and learning capacity from a large amount of data. Transformer-based models have been overwhelmingly preferred (99 times), highlighting their growing dominance in NLP. Large Language Models have also recently been investigated in this area due to their excellent emergent ability.
Evaluation metrics. To evaluate model performance in cross-lingual offensive language detection, macro F1 is the most used metric given class imbalance in datasets, while accuracy is also popular. Some studies also consider precision, recall, weighted F1 score for a detailed assessment. Eronen et al. (2022) propose a new linguistic similarity metric based on the World Atlas of Language Structures (WALS) (Dryer and Haspelmath, 2013) to select optimal transfer languages for automatic hate speech detection.
Source code availability. Out of 67 papers reviewed, 26 (3̃8%) made their source code available, whereas the remainder 41 (62%) do not, hindering transparency and reproducibility.
Language pairs. We present the frequency of language pairs considered in cross-lingual studies in Figure 7. English prevails as the source language, while the most frequent target languages are Spanish, German, and Italian. These languages share typological similarities with English, which likely facilitates their frequent pairing in cross-lingual hate speech detection. This emphasises the impact of data availability and language similarity on the training and application of cross-lingual methods, while also revealing the limitation of relevant cross-lingual research resulting from data scarcity. Many studies have also explored cross-lingual learning by transferring English to Arabic, Hindi and French, as these languages have a relatively large amount of datasets available (see Figure 7). Additionally, there are studies working on the transfer between non-English languages, among which Danish, Indonesian, Portuguese, and Turkish are frequent.
7. Challenges and Future Directions
In this section, we examine key challenges encountered in cross-lingual offensive language detection in three aspects (language, dataset and approach), and then outline potential future directions.
7.1. Language-Related Challenges
7.1.1. Diverse Linguistic Structures
Language diversity encompasses various aspects from basic grammatical rules to complex nuances unique to each language. Since languages within a certain language family may exhibit similarities, cross-lingual transfer works better in more closely related languages and poses challenges for more dissimilar languages (R”̈ottger et al., 2022b). However, this does not necessarily imply a consistency in linguistic features (Eronen et al., 2022). Furthermore, the presence of dialectal forms within a single language amplifies its complexity. For example, the Arabic language is spoken by a wide range of countries across Asia, with complexities arising due to its multiple dialects (Husain and Uzuner, 2022).
7.1.2. Code Mixing
Non-English social media content is often characterised by code-mixing, where users blend languages, use transliterations, or incorporate multiple scripts within a single sentence or conversation, such as English-Hindi (Ibrohim and Budi, 2019a; Hande et al., 2021b) and English-Chinese (Jiang and Zubiaga, 2023). The English-Hindi combinations (also called Hinglish) appear in many cases, because the unique policies on regulating such speech in these regions add layers of complexity and make detection more challenging (Kapoor et al., 2019).
7.1.3. Cultural Variations
Culture is multifaceted and complex. The perception of offensive language can differ significantly across languages and cultures (Nozza, 2021). What is considered non-offensive in one culture or language might be misinterpreted as signals of offense in another, leading to detection discrepancies. Even within a single language, there can be cultural diversity. For example, categorising English as a representative of “western cultural background” might ignore the differences between American and British cultures (Zhou et al., 2023). Such cultural variations can hinder the effective transferability of language models (R”̈ottger et al., 2022b). Moreover, hateful expressions often employ figurative language, rhetorical figures and idioms, which are also language-dependent. This requires models to interpret underlying intent rather than relying solely on literal meanings (Pamungkas et al., 2021a).
7.2. Dataset-Related Challenges
7.2.1. Limited Labelled Datasets
Low-resource languages often do not have the attention that major languages receive, leading to limited data availability (Burtenshaw and Kestemont, 2021) and limited studies (see Figures 3 and 7). Additionally, creating a high-quality labelled dataset requires extensive manual effort, which can potentially lead to high annotation and time costs as well as be harmful to annotators (R”̈ottger et al., 2022a). Annotating offensive language is even more challenging due to the need for cultural and linguistic knowledge. Ethical and privacy concerns also arise when dealing with real-world data from social media platforms or other online sources (Jiang et al., 2022), because there are potential risks of exposing sensitive information or inadvertently promoting offensive language when studying it or trying to combat it.
7.2.2. Inconsistent Definition of Offensive Language
The diversity of hate speech encompassing from misogyny and racism to other forms of discrimination often leads to discrepancies in research (Nozza, 2021). This has a bigger impact on low-resource languages (Bigoulaeva et al., 2021, 2022, 2023), and introduce ambiguities that challenge the task, especially in multilingual scenarios (Vitiugin et al., 2021). The lack of definition consistency across datasets in turn limits cross-lingual research, because many datasets in either source or target languages may be incompatible (Bigoulaeva et al., 2022).
7.2.3. Inconsistent Annotations
Offensive language datasets can exhibit systematic gaps and biases in their annotations (R”̈ottger et al., 2022b). Different labelled datasets often utilise distinct annotation frameworks or strategies (Fortuna and Nunes, 2018). The subjective nature of offensive language leads to tentatively disputed annotations among human annotators, where hateful content could be interpreted differently depending on annotators’ understanding (Das et al., 2022). Annotating these contents is easily influenced by individual sociodemographic and cultural backgrounds, often resulting in low inter-annotator agreement (Markov et al., 2021). Additionally, there is an “annotation’s dilemma” (Aluru et al., 2021), whereby ambiguous instances might be annotated incorrectly by annotators but predicted accurately by the model.
7.2.4. Dataset Imbalance
Given that the majority of social media content is non-hateful, labels are typically skewed towards the non-hate label (Song et al., 2021; Bigoulaeva et al., 2022, 2023). This skewness can lead to training issues, especially when working with small training corpora. In addition, when considering multilingual or merged datasets, inter-language imbalance is another challenge (Deshpande et al., 2022). Such datasets often display significant differences in the number of examples available for each language. This imbalance can bias models towards the semantic tendencies of languages that are overrepresented.
7.2.5. Dataset Bias
Biases such as topic, authorship, political affiliation, social media platform, and collection period can be one of the major issues and lead to potential generalisation challenges for cross-lingual models. Among them, topic bias has been taken seriously in many works (R”̈ottger et al., 2022a; Arango et al., 2022; Vitiugin et al., 2021; R”̈ottger et al., 2022b). While some datasets include general offensive language or hate speech, others might concentrate on specific issues such as immigrants, misogyny, politics, or religion. Such topic-specific biases across datasets can harm model performance in cross-lingual classification (Pamungkas et al., 2021a; Stappen et al., 2020). Besides, the temporal aspect of data collection, affected by events in different periods, can further introduce bias, leading to datasets with varied topical foci (Stappen et al., 2020).
7.2.6. Data Source Obsolescence
As online language rapidly transforms, especially influenced by moderation and real-world events, datasets can quickly become obsolete (Vitiugin et al., 2021; Alkhalifa et al., 2023). This obsolescence is further exacerbated by the emergence of new slang, metaphors, and colloquialisms that vary across languages and regions. For instance, terms that normally have neutral meanings (such as donkey) may be used offensively in sexist text (Jiang and Zubiaga, 2023). The constrained and informal style of tweets, which are more like oral expressions than written language, complicates the preprocessing steps and leads to erroneous predictions (Pamungkas et al., 2020; Aluru et al., 2021). Emojis, associated with various forms of online harassment, further pose unique challenges (Corazza et al., 2020). Moreover, the domain and target of hate speech can shift significantly over time (Florio et al., 2020). Real-world events, from local incidents to global crises like the COVID-19 pandemic, can give rise to new domains and terms of hate speech (Montariol et al., 2022). Then, it becomes increasingly challenging to create and continually update datasets customised for every possible language and domain, leading to frequent low-resource issues.
7.3. Approach-Related Challenges
7.3.1. Limited Ability of Multilingual Pre-Trained Language Models
Cross-lingual models, especially multilingual PLMs, often face challenges related to their generalisation abilities (Hahn et al., 2021; Pelicon et al., 2021b). The generalisability of transformer-based multilingual PLMs can be inconsistent, especially for typologically diverse languages, because they might be pre-trained with a highly focused set of languages or some languages with insufficient training samples (Deshpande et al., 2022). This can lead to bias and instability in model performance, depending on the model architectures or topical focus in datasets (Mozafari et al., 2022; Pamungkas et al., 2021a). Model overfitting across target languages is another important issue. Although multilingual models are less prone to overfitting on dataset-specific features than monolingual models, they will achieve poorer performance than monolingual ones in higher-resource settings, and they may require very different calibrations and adaptations across languages (R”̈ottger et al., 2022a, b).
7.3.2. Cross-lingual Transfer Performance
CLTL has emerged as a promising solution and has indeed shown effectiveness when addressing data scarcity (Bigoulaeva et al., 2022; Sohn and Lee, 2019). However, offensive language is deeply rooted in the specificity and diversity of language and culture, and its complexity poses significant obstacles to cross-lingual transfer, especially in zero-shot settings (Nozza, 2021; Montariol et al., 2022). Zero-shot transfer approaches to multilingual training often suffer from performance deficiencies when compared to models trained on actual target language data (Pelicon et al., 2021b).
7.3.3. Limitations of Machine Translation
Machine translation can help augment datasets and alleviate data scarcity issues. A prevalent practice is to utilise Google Translate API.181818https://translate.google.com (Aluru et al., 2021; Pamungkas et al., 2021a; Jiang and Zubiaga, 2021; zahra El-Alami et al., 2022) However, the effectiveness of cross-lingual detection models is directly linked to the quality and accuracy of the machine translation. True semantics of the target text may change during the translation process. Machine translation may diminish the toxicity degree and thereby reduce the perception of offensive content, especially in context-sensitive situations (Das et al., 2022). Therefore, models that rely on such translations can experience performance degradation, especially when predicting instances with semantic shifts. However, despite translation errors and uncertainties, they can still be valuable supplementary inputs for text classification (Sohn and Lee, 2019).
7.3.4. Poor Model Interpretability
While advanced deep learning-based models have shown good performance in detecting hate speech, they often operate as “black boxes” with a lack of transparency in the decision-making process (Aluru et al., 2021). This opacity makes it hard for humans to understand the underlying reasons for the model performance in a particular language or scenario, and analyse the model errors (Deshpande et al., 2022; Vitiugin et al., 2021). To bridge this interpretability gap, some tools are utilised to explain what models are doing, such as Local Interpretable Model-agnostic Explanations (LIME) (Ribeiro et al., 2016) and SHapley Additive exPlanations (SHAP) (Lundberg and Lee, 2017), and theoretical approaches from cognitive linguistics are also applied, e.g. frame semantics (Vitiugin et al., 2021). A human-in-the-loop paradigm can also help by integrating human expertise into the model’s learning process, but it can be difficult and uncertain for individuals to provide effective feedback to enhance the model performance (Vitiugin et al., 2021).
7.4. Future Directions
This section outlines promising research directions in this critical research area.
7.4.1. Dataset Creation
There is a pressing need to collect and annotate comprehensive and balanced datasets, especially in low-resource languages (Husain and Uzuner, 2022; Arango Monnar et al., 2022; Gaikwad et al., 2021; Burtenshaw and Kestemont, 2021). While zero-shot cross-lingual transfer has its limitations, even a small amount of target data can substantially enhance model fine-tuning (R”̈ottger et al., 2022a). To avoid dataset bias, such datasets can span multiple languages, dialects, cultural contexts, and diverse types of offensive language (Ousidhoum et al., 2019). More diverse strategies for data collection, such as sampling from various platforms and user demographics, can yield richer and more annotation-worthy data (R”̈ottger et al., 2022a). Additionally, the creation of domain-specific evaluation datasets, such as XHATE-999 (Glavaš et al., 2020) and Multilingual HateCheck (MHC) (R”̈ottger et al., 2022b), provides a multifaceted and precise perspective for assessing cross-lingual detection.
7.4.2. Data Annotation
Refining data annotation strategies is instrumental. Instead of large annotating tasks, annotation can be done incrementally, ensuring efficiency and encouraging diverse data collection (R”̈ottger et al., 2022a). In addition, semi-supervised learning methods can efficiently use data and annotator resources and mitigate data scarcity issues. For example, deliberate selection strategies, such as active learning, can enhance annotation performance with fewer target-language entries, mainly annotating data points that the model finds most ambiguous (R”̈ottger et al., 2022a). To further address ambiguity and sensitivity in offensive language, incorporating feedback from human experts can be invaluable (Vitiugin et al., 2021). It would also be better for annotators to follow ethical guidelines and receive appropriate training and support, avoiding potential harm to them.
7.4.3. Integration of Additional Features
Integrating additional features can also be an effective strategy. Typographic features, including capitals, punctuation, and emojis, offer a language-agnostic and domain-agnostic perspective to offensive language detection. Emojis, associated with emotional expressions, have been highlighted for their ability to help identify online hate due to their shared meanings across languages (Kumar and Sachdeva, 2022; Markov et al., 2021; Corazza et al., 2020). Stylometric features are also robust indicators of offensive language, as the writing style of toxic content can be correlated to social media users (Markov et al., 2021). In addition, domain-specific features are a critical aspect in additional features to assist the knowledge transfer process, especially when addressing linguistic nuances like metaphors, metonymy, and informal expressions. The multilingual lexicon, HurtLex, has been highlighted for its effectiveness in multiple studies (Pamungkas et al., 2020, 2021a; Pamungkas and Patti, 2019; Jiang and Zubiaga, 2021). Since general-purpose multilingual embeddings may not effectively capture some hate-specific patterns, some researchers also utilise hateful word pairs to construct domain-aware or semantic-based representations across languages (Jiang and Zubiaga, 2023; Hahn et al., 2021; Arango et al., 2021). Furthermore, infusing cultural features underscores the significance of language-specific knowledge for geographically sensitive tasks. The creation of culturally-aware models, informed by multidisciplinary studies involving anthropologists and sociologists, can provide a richer understanding of cultural dimensions in hateful content (Zhou et al., 2023). Such work can produce different forms of cultural features, such as switching pattern matrix (Bansal et al., 2020), bilingual language pairs (Jiang and Zubiaga, 2023), cross-cultural similarities (Arango et al., 2021), and social dynamics among users (Nozza, 2021).
7.4.4. Multilingual Pre-Trained Language Models
Multilingual PLMs can be applied to low-resource target languages without further training, thus bridging the data availability gap between high- and low-resource languages. However, there are significant computational costs during training and fine-tuning due to their large number of parameters (Pikuliak et al., 2021). Future research is geared towards optimising these multilingual PLMs for enhanced efficiency and interpretability. Additionally, two strategies are emerging to improve the generalisability and scalability of multilingual PLMs so that they can scale to handle large datasets or multiple languages simultaneously (Arango et al., 2022). The first emphasises pre-training models using data from relevant sources (such as social platforms and hateful domains), as demonstrated by XLM-T (Montariol et al., 2022) and AbuseXLMR (Gupta et al., 2022). The second strategy focuses on models based on specific low-resource languages or those from similar language families, as seen with MuRIL for Indic languages (Das et al., 2022) and AraBERT for dialects (Husain and Uzuner, 2022).
7.4.5. Cross-lingual Training Strategies
Researchers are also delving into a number of innovative cross-lingual training strategies. Adaptive training techniques can dynamically adjust to the specificities of different languages and dialects, as well as bridge the gap between offensive language detection and PLMs. By leveraging meta-learning or few-shot learning, models can be trained to quickly adapt to new low-resource languages with few labelled data (Mozafari et al., 2022). Furthermore, collaborative multi-task learning can make models more generalisable in cultural variations and resistant to overfitting by training them and sharing representations on multiple auxiliary tasks simultaneously across different languages or similar NLP tasks (Montariol et al., 2022). The use of adversarial training, where models are trained against adversarial examples, can also enhance their robustness and generalisation across languages. This is especially important for offensive language, as attackers often use subtle language tricks to obfuscate and bypass detection (Shi et al., 2022; Jin et al., 2020; Ryzhova et al., 2022). Moreover, cascade learning can greatly bolster model performance and enhance out-of-domain generalisability by initially training models on source data or more diverse datasets and subsequently fine-tuning the results by simultaneously or progressively incorporating target language data (R”̈ottger et al., 2022a; Pelicon et al., 2021b). Besides, Vitiugin et al. (2021) highlight that human feedback can be of great value to help detect subtleties of abuse and phrases during model training, Ranasinghe and Zampieri (2021c) explore language-specific preprocessing like segmentation in morphologically rich languages such as Arabic and Turkish, and multi-source training, with augmented datasets, is also able to capture the diversities and commonalities of linguistic patterns across languages (Hande et al., 2021a; Markov et al., 2021; Chiril et al., 2019).
7.4.6. Application of Large Language Models
The emergence of Large Language Models (LLMs), such as GPT-3 (Brown et al., 2020), GPT-4 (OpenAI, 2023), Fine-tuned LAnguage Net (FLAN) (Wei et al., 2021) and Large Language Model Meta AI (LlaMA) (Touvron et al., 2023), has opened up new horizons for cross-lingual offensive language detection, thanks to their remarkable performance (Zhao et al., 2023). Instruction fine-tuned LLMs with prompting have shown promise in detecting offensive language across both mono and multilingual contents without the need for language-specific fine-tuning (Plaza-del arco et al., 2023). The efficacy of LLMs can be further enhanced through well-designed, task-specific prompts (Chen et al., 2022; Han and Tang, 2022; Roy et al., 2023). By incorporating informative task descriptions and input-label demonstrations, these prompts can guide LLMs to achieve superior detection. Furthermore, the inherent multi-tasking capabilities of LLMs allow them to excel not only in offensive language detection but also in some related NLP tasks (Zhang et al., 2023). To address the challenge of data scarcity, LLMs offer a potential solution by synthetically generating hateful texts, especially for languages with limited resources (Hartvigsen et al., 2022). Beyond detection, LLMs can also produce explanations for predictions, enhancing both model performance and interpretability (Wang et al., 2023). While the potential of LLMs in this domain is evident, current research is still in its infancy.
8. Conclusion
This survey provides a detailed exploration of CLTL techniques in detecting offensive language online, Our survey is the first to specifically focus on cross-lingual detection, analysing a collection of 67 papers. We summarise the characteristics of multilingual offensive language datasets and cross-lingual resources. We further emphasise the diversity and complexity of CLTL techniques, highlighting the varying levels of transfer –instance, feature, and parameter– and popular strategies to improve performance. The survey also reveals several ongoing challenges in cross-lingual offensive language detection, such as linguistic diversity, dataset scarcity, and model generalisability. We also point out potential research opportunities, emphasising the need for new datasets across low-resource languages, as well as more robust, generalisable, and ethically sound CLTL approaches. Our survey highlights the importance of advancing CLTL approaches, which can serve as both a key reference for current practices and a guide for future research in this evolving field.
While providing a clear focus and an in-depth discussion on cross-lingual resources and methods, our survey does not cover offensive language datasets, resources and methods outside this scope.
Acknowledgements.
Aiqi Jiang is funded by China Scholarship Council (CSC Funding, No. 201908510140). Arkaitz Zubiaga acknowledges support from the European Union and UK Research and Innovation under Grant No. 101073351 as part of Marie Skłodowska-Curie Actions (MSCA Hybrid Intelligence to monitor, promote, and analyze transformations in good democracy practices).References
- (1)
- Akhter et al. (2020) Muhammad Pervez Akhter, Zheng Jiangbin, Irfan Raza Naqvi, Mohammed Abdelmajeed, and Muhammad Tariq Sadiq. 2020. Automatic detection of offensive language for urdu and roman urdu. IEEE Access 8 (2020), 91213–91226.
- Al-Hassan and Al-Dossari (2019) Areej Al-Hassan and Hmood Al-Dossari. 2019. Detection of hate speech in social networks: a survey on multilingual corpus. In Proc. of COMIT, Vol. 10. 10–5121.
- Alakrot et al. (2018) Azalden Alakrot, Liam Murray, and Nikola S Nikolov. 2018. Dataset construction for the detection of anti-social behaviour in online communication in Arabic. Procedia Computer Science 142 (2018), 174–181.
- Alfina et al. (2017) Ika Alfina, Rio Mulia, Mohamad Ivan Fanany, and Yudo Ekanata. 2017. Hate speech detection in the Indonesian language: A dataset and preliminary study. In Proc. of ICACSIS. 233–238.
- Alkhalifa et al. (2023) Rabab Alkhalifa, Elena Kochkina, and Arkaitz Zubiaga. 2023. Building for tomorrow: Assessing the temporal persistence of text classifiers. Information Processing & Management 60, 2 (2023), 103200.
- Alkomah and Ma (2022) Fatimah Alkomah and Xiaogang Ma. 2022. A literature review of textual hate speech detection methods and datasets. Information 13, 6 (2022), 273.
- Aluru et al. (2020) Sai Saketh Aluru, Binny Mathew, Punyajoy Saha, and Animesh Mukherjee. 2020. Deep learning models for multilingual hate speech detection. arXiv preprint arXiv:2004.06465 (2020).
- Aluru et al. (2021) Sai Saketh Aluru, Binny Mathew, Punyajoy Saha, and Animesh Mukherjee. 2021. A Deep Dive into Multilingual Hate Speech Classification. In Proc. of ECML PKDD, Yuxiao Dong, Georgiana Ifrim, Dunja Mladenić, Craig Saunders, and Sofie Van Hoecke (Eds.). Springer, Cham, 423–439.
- Amjad et al. (2022) Maaz Amjad, Alisa Zhila, Grigori Sidorov, Andrey Labunets, Sabur Butt, Hamza Imam Amjad, Oxana Vitman, and Alexander Gelbukh. 2022. UrduThreat@FIRE2021: Shared Track on Abusive Threat Identification in Urdu. In Proc. of FIRE (Virtual Event, India). ACM, New York, NY, USA, 9–11.
- Arango et al. (2019) Aymé Arango, Jorge Pérez, and Barbara Poblete. 2019. Hate Speech Detection is Not as Easy as You May Think: A Closer Look at Model Validation. In Proc. of ACM SIGIR (Paris, France). ACM, New York, NY, USA, 45–54.
- Arango et al. (2021) Aymé Arango, Jorge Pérez, and Barbara Poblete. 2021. Cross-lingual hate speech detection based on multilingual domain-specific word embeddings. arXiv preprint arXiv:2104.14728 (2021).
- Arango et al. (2022) Aymé Arango, Jorge Pérez, and Barbara Poblete. 2022. Hate speech detection is not as easy as you may think: A closer look at model validation (extended version). Information Systems 105 (2022), 101584.
- Arango Monnar et al. (2022) Ayme Arango Monnar, Jorge Perez, Barbara Poblete, Magdalena Saldaña, and Valentina Proust. 2022. Resources for Multilingual Hate Speech Detection. In Proc. of the WOAH Workshop. ACL, Seattle, Washington (Hybrid), 122–130.
- Arata (2019) Masaki Arata. 2019. Study on change of detection accuracy over time in cyberbullying detection. Ph. D. Dissertation. Master’s thesis, Kitami Institute of Technology, Department of Computer Science.
- Artetxe and Schwenk (2019) Mikel Artetxe and Holger Schwenk. 2019. Massively Multilingual Sentence Embeddings for Zero-Shot Cross-Lingual Transfer and Beyond. Transactions of the ACL 7 (2019), 597–610.
- Awal et al. (2023) Md Rabiul Awal, Roy Ka-Wei Lee, Eshaan Tanwar, Tanmay Garg, and Tanmoy Chakraborty. 2023. Model-Agnostic Meta-Learning for Multilingual Hate Speech Detection. IEEE Trans. on Computational Social Systems (2023), 1–10.
- Ayo et al. (2020) Femi Emmanuel Ayo, Olusegun Folorunso, Friday Thomas Ibharalu, and Idowu Ademola Osinuga. 2020. Machine learning techniques for hate speech classification of twitter data: State-of-the-art, future challenges and research directions. Computer Science Review 38 (2020), 100311.
- Bansal et al. (2020) Srijan Bansal, Vishal Garimella, Ayush Suhane, Jasabanta Patro, and Animesh Mukherjee. 2020. Code-Switching Patterns Can Be an Effective Route to Improve Performance of Downstream NLP Applications: A Case Study of Humour, Sarcasm and Hate Speech Detection. In Proc. of ACL. ACL, Online, 1018–1023.
- Basile et al. (2019) Valerio Basile, Cristina Bosco, Elisabetta Fersini, Debora Nozza, Viviana Patti, Francisco Manuel Rangel Pardo, Paolo Rosso, and Manuela Sanguinetti. 2019. SemEval-2019 Task 5: Multilingual Detection of Hate Speech Against Immigrants and Women in Twitter. In Proc. of the SemEval Workshop. ACL, Minneapolis, Minnesota, USA, 54–63.
- Bhattacharya et al. (2020) Shiladitya Bhattacharya, Siddharth Singh, Ritesh Kumar, Akanksha Bansal, Akash Bhagat, Yogesh Dawer, Bornini Lahiri, and Atul Kr Ojha. 2020. Developing a multilingual annotated corpus of misogyny and aggression. arXiv preprint arXiv:2003.07428 (2020).
- Bigoulaeva et al. (2021) Irina Bigoulaeva, Viktor Hangya, and Alexander Fraser. 2021. Cross-Lingual Transfer Learning for Hate Speech Detection. In Proc. of the LT-EDI Workshop. ACL, Kyiv, 15–25.
- Bigoulaeva et al. (2022) Irina Bigoulaeva, Viktor Hangya, Iryna Gurevych, and Alexander Fraser. 2022. Addressing the Challenges of Cross-Lingual Hate Speech Detection. arXiv preprint arXiv:2201.05922 (2022).
- Bigoulaeva et al. (2023) Irina Bigoulaeva, Viktor Hangya, Iryna Gurevych, and Alexander Fraser. 2023. Label modification and bootstrapping for zero-shot cross-lingual hate speech detection. Language Resources and Evaluation (2023), 1–32.
- Bohra et al. (2018) Aditya Bohra, Deepanshu Vijay, Vinay Singh, Syed Sarfaraz Akhtar, and Manish Shrivastava. 2018. A Dataset of Hindi-English Code-Mixed Social Media Text for Hate Speech Detection. In Proc. of PEOPLES Workshop. ACL, New Orleans, Louisiana, USA, 36–41.
- Bosco et al. (2018) Cristina Bosco, Dell’Orletta Felice, Fabio Poletto, Manuela Sanguinetti, Tesconi Maurizio, et al. 2018. Overview of the evalita 2018 hate speech detection task. In Proc. of CEUR Workshop, Vol. 2263. CEUR, 1–9.
- Boulouard et al. (2022) Zakaria Boulouard, Mariya Ouaissa, Mariyam Ouaissa, Moez Krichen, Mutiq Almutiq, and Karim Gasmi. 2022. Detecting Hateful and Offensive Speech in Arabic Social Media Using Transfer Learning. Applied Sciences 12, 24 (2022).
- Bretschneider and Peters (2017) Uwe Bretschneider and Ralf Peters. 2017. Detecting offensive statements towards foreigners in social media. (2017).
- Brown et al. (2020) Tom Brown, Benjamin Mann, Nick Ryder, Melanie Subbiah, Jared D Kaplan, Prafulla Dhariwal, Arvind Neelakantan, Pranav Shyam, Girish Sastry, Amanda Askell, et al. 2020. Language models are few-shot learners. Proc. of NeurIPS 33 (2020), 1877–1901.
- Burtenshaw and Kestemont (2021) Ben Burtenshaw and Mike Kestemont. 2021. A Dutch Dataset for Cross-lingual Multilabel Toxicity Detection. In Proc. of the BUCC Workshop. INCOMA Ltd., Online (Virtual Mode), 75–79.
- Chakravarthi et al. (2022) Bharathi Raja Chakravarthi, Ruba Priyadharshini, Vigneshwaran Muralidaran, Navya Jose, Shardul Suryawanshi, Elizabeth Sherly, and John P McCrae. 2022. Dravidiancodemix: Sentiment analysis and offensive language identification dataset for dravidian languages in code-mixed text. Language Resources and Evaluation 56, 3 (2022), 765–806.
- Charitidis et al. (2020) Polychronis Charitidis, Stavros Doropoulos, Stavros Vologiannidis, Ioannis Papastergiou, and Sophia Karakeva. 2020. Towards countering hate speech against journalists on social media. Online Social Networks and Media 17 (2020), 100071.
- Chen et al. (2022) Yulong Chen, Yang Liu, Li Dong, Shuohang Wang, Chenguang Zhu, Michael Zeng, and Yue Zhang. 2022. AdaPrompt: Adaptive Model Training for Prompt-based NLP. In Findings of EMNLP. ACL, Abu Dhabi, United Arab Emirates, 6057–6068.
- Chhabra and Vishwakarma (2023) Anusha Chhabra and Dinesh Kumar Vishwakarma. 2023. A literature survey on multimodal and multilingual automatic hate speech identification. Multimedia Systems (2023), 1–28.
- Chiril et al. (2019) Patricia Chiril, Farah Benamara Zitoune, Véronique Moriceau, Marlène Coulomb-Gully, and Abhishek Kumar. 2019. Multilingual and Multitarget Hate Speech Detection in Tweets. In Proc. of TALN-RECITAL. ATALA, Toulouse, France, 351–360.
- Chung et al. (2019) Yi-Ling Chung, Elizaveta Kuzmenko, Serra Sinem Tekiroglu, and Marco Guerini. 2019. CONAN - COunter NArratives through Nichesourcing: a Multilingual Dataset of Responses to Fight Online Hate Speech. In Proc. of ACL. ACL, Florence, Italy, 2819–2829.
- Çöltekin (2020) Çağrı Çöltekin. 2020. A corpus of Turkish offensive language on social media. In Proc. of LREC. 6174–6184.
- Conneau et al. (2020) Alexis Conneau, Kartikay Khandelwal, Naman Goyal, Vishrav Chaudhary, Guillaume Wenzek, Francisco Guzmán, Edouard Grave, Myle Ott, Luke Zettlemoyer, and Veselin Stoyanov. 2020. Unsupervised Cross-lingual Representation Learning at Scale. In Proc. of ACL. ACL, Online, 8440–8451.
- Corazza et al. (2020) Michele Corazza, Stefano Menini, Elena Cabrio, Sara Tonelli, and Serena Villata. 2020. Hybrid Emoji-Based Masked Language Models for Zero-Shot Abusive Language Detection. In Findings of EMNLP. ACL, Online, 943–949.
- Das et al. (2022) Mithun Das, Somnath Banerjee, and Animesh Mukherjee. 2022. Data Bootstrapping Approaches to Improve Low Resource Abusive Language Detection for Indic Languages. In Proc. of ACM HT (Barcelona, Spain). ACM, New York, NY, USA, 32–42.
- Davidson et al. (2017) Thomas Davidson, Dana Warmsley, Michael Macy, and Ingmar Weber. 2017. Automated Hate Speech Detection and the Problem of Offensive Language. In Proc. of ICWSM (Montreal, Canada). 512–515.
- de Gibert et al. (2018) Ona de Gibert, Naiara Perez, Aitor García-Pablos, and Montse Cuadros. 2018. Hate Speech Dataset from a White Supremacy Forum. In Proc. of the ALW Workshop. ACL, Brussels, Belgium, 11–20.
- De la Peña Sarracén and Rosso (2022) Gretel Liz De la Peña Sarracén and Paolo Rosso. 2022. Unsupervised Embeddings with Graph Auto-Encoders for Multi-domain and Multilingual Hate Speech Detection. In Proc. of LREC. ELRA, Marseille, France, 2196–2204.
- de Pelle and Moreira (2017) Rogers de Pelle and Viviane Moreira. 2017. Offensive Comments in the Brazilian Web: a dataset and baseline results. In Anais do VI Brazilian Workshop on Social Network Analysis and Mining (São Paulo). SBC, Porto Alegre, RS, Brasil.
- Demilie and Salau (2022) Wubetu Barud Demilie and Ayodeji Olalekan Salau. 2022. Detection of fake news and hate speech for Ethiopian languages: a systematic review of the approaches. Journal of Big Data 9, 1 (2022), 66.
- Deng et al. (2022) Jiawen Deng, Jingyan Zhou, Hao Sun, Chujie Zheng, Fei Mi, Helen Meng, and Minlie Huang. 2022. COLD: A Benchmark for Chinese Offensive Language Detection. In Proc. of EMNLP. ACL, Abu Dhabi, United Arab Emirates, 11580–11599.
- Deshpande et al. (2022) Neha Deshpande, Nicholas Farris, and Vidhur Kumar. 2022. Highly generalizable models for multilingual hate speech detection. arXiv preprint arXiv:2201.11294 (2022).
- Devlin et al. (2019) Jacob Devlin, Ming-Wei Chang, Kenton Lee, and Kristina Toutanova. 2019. BERT: Pre-training of Deep Bidirectional Transformers for Language Understanding. In Proc. of NAACL-HLT. ACL, Minneapolis, Minnesota, 4171–4186.
- Devyatkin et al. (2017) Dmitry Devyatkin, Ivan Smirnov, Margarita Ananyeva, Maria Kobozeva, Andrey Chepovskiy, and Fyodor Solovyev. 2017. Exploring linguistic features for extremist texts detection (on the material of Russian-speaking illegal texts). In Proc. of IEEE ISI. 188–190.
- Dowlagar and Mamidi (2021) Suman Dowlagar and Radhika Mamidi. 2021. A survey of recent neural network models on code-mixed indian hate speech data. In Proc. of FIRE. 67–74.
- Dryer and Haspelmath (2013) Matthew S. Dryer and Martin Haspelmath (Eds.). 2013. WALS Online (v2020.3). Zenodo. https://doi.org/10.5281/zenodo.7385533
- Dybala et al. (2010) Michal Ptaszynski1 Pawel Dybala, Tatsuaki Matsuba2 Fumito Masui, Rafal Rzepka, and Kenji Araki. 2010. Machine Learning and Affect Analysis against Cyber-Bullying. In Proc. of the Linguistic and Cognitive Approaches to Dialog Agents Symposium.
- Elouali et al. (2020) Aya Elouali, Zakaria Elberrichi, and Nadia Elouali. 2020. Hate Speech Detection on Multilingual Twitter Using Convolutional Neural Networks. Revue d’Intelligence Artificielle 34, 1 (2020), 81–88.
- ElSherief et al. (2018) Mai ElSherief, Shirin Nilizadeh, Dana Nguyen, Giovanni Vigna, and Elizabeth Belding. 2018. Peer to peer hate: Hate speech instigators and their targets. In Proc. of ICWSM, Vol. 12.
- Eronen et al. (2022) Juuso Eronen, Michal Ptaszynski, Fumito Masui, Masaki Arata, Gniewosz Leliwa, and Michal Wroczynski. 2022. Transfer language selection for zero-shot cross-lingual abusive language detection. Information Processing & Management 59, 4 (2022), 102981.
- Feng et al. (2022) Fangxiaoyu Feng, Yinfei Yang, Daniel Cer, Naveen Arivazhagan, and Wei Wang. 2022. Language-agnostic BERT Sentence Embedding. In Proc. of ACL. ACL, Dublin, Ireland, 878–891.
- Ferreira et al. (2016) Daniel C. Ferreira, André F. T. Martins, and Mariana S. C. Almeida. 2016. Jointly Learning to Embed and Predict with Multiple Languages. In Proc. of ACL). ACL, Berlin, Germany, 2019–2028.
- Fersini et al. (2018a) Elisabetta Fersini, Debora Nozza, and Paolo Rosso. 2018a. Overview of the Evalita 2018 Task on Automatic Misogyny Identification (AMI). In EVALITA@CLiC-it.
- Fersini et al. (2018b) Elisabetta Fersini, Paolo Rosso, and Mary E. Anzovino. 2018b. Overview of the Task on Automatic Misogyny Identification at IberEval 2018. In IberEval@SEPLN.
- Finn et al. (2017) Chelsea Finn, Pieter Abbeel, and Sergey Levine. 2017. Model-agnostic meta-learning for fast adaptation of deep networks. In Proc. of ICML. PMLR, 1126–1135.
- Firmino et al. (2021) Anderson Almeida Firmino, Cláudio Souza de Baptista, and Anselmo Cardoso de Paiva. 2021. Using Cross Lingual Learning for Detecting Hate Speech in Portuguese. In Database and Expert Systems Applications, Christine Strauss, Gabriele Kotsis, A. Min Tjoa, and Ismail Khalil (Eds.). Springer, Cham, 170–175.
- Florio et al. (2020) Komal Florio, Valerio Basile, Marco Polignano, Pierpaolo Basile, and Viviana Patti. 2020. Time of your hate: The challenge of time in hate speech detection on social media. Applied Sciences 10, 12 (2020), 4180.
- for Peace ([n. d.]) Rights for Peace. [n. d.]. What is Hate Speech? https://www.rightsforpeace.org/hate-speech. Accessed: 2023-10-24.
- Fortuna (2017) Paula Fortuna. 2017. Automatic detection of hate speech in text: an overview of the topic and dataset annotation with hierarchical classes.
- Fortuna and Nunes (2018) Paula Fortuna and Sérgio Nunes. 2018. A Survey on Automatic Detection of Hate Speech in Text. 51, 4, Article 85 (jul 2018), 30 pages.
- Fortuna et al. (2019) Paula Fortuna, João Rocha da Silva, Juan Soler-Company, Leo Wanner, and Sérgio Nunes. 2019. A Hierarchically-Labeled Portuguese Hate Speech Dataset. In Proc. of the ALW Workshop. ACL, Florence, Italy, 94–104.
- Founta et al. (2018) Antigoni Founta, Constantinos Djouvas, Despoina Chatzakou, Ilias Leontiadis, Jeremy Blackburn, Gianluca Stringhini, Athena Vakali, Michael Sirivianos, and Nicolas Kourtellis. 2018. Large scale crowdsourcing and characterization of twitter abusive behavior. In Proc. of ICWSM, Vol. 12.
- Gaikwad et al. (2021) Saurabh Sampatrao Gaikwad, Tharindu Ranasinghe, Marcos Zampieri, and Christopher Homan. 2021. Cross-lingual Offensive Language Identification for Low Resource Languages: The Case of Marathi. In Proc. of RANLP. INCOMA Ltd., Held Online, 437–443.
- Gao and Huang (2017) Lei Gao and Ruihong Huang. 2017. Detecting online hate speech using context aware models. arXiv preprint arXiv:1710.07395 (2017).
- Garg et al. (2023) Tanmay Garg, Sarah Masud, Tharun Suresh, and Tanmoy Chakraborty. 2023. Handling Bias in Toxic Speech Detection: A Survey. 55, 13s, Article 264 (jul 2023), 32 pages.
- Glavaš et al. (2020) Goran Glavaš, Mladen Karan, and Ivan Vulić. 2020. XHate-999: Analyzing and Detecting Abusive Language Across Domains and Languages. In Proc. of COLING. ICCL, Barcelona, Spain (Online), 6350–6365.
- Golbeck et al. (2017) Jennifer Golbeck, Zahra Ashktorab, Rashad O. Banjo, Alexandra Berlinger, Siddharth Bhagwan, Cody Buntain, Paul Cheakalos, Alicia A. Geller, Quint Gergory, Rajesh Kumar Gnanasekaran, Raja Rajan Gunasekaran, Kelly M. Hoffman, Jenny Hottle, Vichita Jienjitlert, Shivika Khare, Ryan Lau, Marianna J. Martindale, Shalmali Naik, Heather L. Nixon, Piyush Ramachandran, Kristine M. Rogers, Lisa Rogers, Meghna Sardana Sarin, Gaurav Shahane, Jayanee Thanki, Priyanka Vengataraman, Zijian Wan, and Derek Michael Wu. 2017. A Large Labeled Corpus for Online Harassment Research. In Proc. of ACM WebSci (Troy, New York, USA). ACM, New York, NY, USA, 229–233.
- Grave et al. (2018) Edouard Grave, Piotr Bojanowski, Prakhar Gupta, Armand Joulin, and Tomas Mikolov. 2018. Learning Word Vectors for 157 Languages. In Proc. of LREC.
- Gupta et al. (2022) Vikram Gupta, Sumegh Roychowdhury, Mithun Das, Somnath Banerjee, Punyajoy Saha, Binny Mathew, Animesh Mukherjee, et al. 2022. Multilingual Abusive Comment Detection at Scale for Indic Languages. Proc. of NeurIPS 35 (2022), 26176–26191.
- Hahn et al. (2021) Vanessa Hahn, Dana Ruiter, Thomas Kleinbauer, and Dietrich Klakow. 2021. Modeling Profanity and Hate Speech in Social Media with Semantic Subspaces. In Proc. of the WOAH Workshop. ACL, Online, 6–16.
- Han and Tang (2022) Lawrence Han and Hao Tang. 2022. Designing of Prompts for Hate Speech Recognition with In-Context Learning. In Proc. of CSCI. IEEE, 319–320.
- Hande et al. (2021a) Adeep Hande, Siddhanth U Hegde, Ruba Priyadharshini, Rahul Ponnusamy, Prasanna Kumar Kumaresan, Sajeetha Thavareesan, and Bharathi Raja Chakravarthi. 2021a. Benchmarking multi-task learning for sentiment analysis and offensive language identification in under-resourced dravidian languages. arXiv preprint arXiv:2108.03867 (2021).
- Hande et al. (2020) Adeep Hande, Ruba Priyadharshini, and Bharathi Raja Chakravarthi. 2020. KanCMD: Kannada CodeMixed Dataset for Sentiment Analysis and Offensive Language Detection. In Proc. of the PEOPLES Workshop. ACL, Barcelona, Spain (Online), 54–63.
- Hande et al. (2021b) Adeep Hande, Karthik Puranik, Konthala Yasaswini, Ruba Priyadharshini, Sajeetha Thavareesan, Anbukkarasi Sampath, Kogilavani Shanmugavadivel, Durairaj Thenmozhi, and Bharathi Raja Chakravarthi. 2021b. Offensive language identification in low-resourced code-mixed dravidian languages using pseudo-labeling. arXiv preprint arXiv:2108.12177 (2021).
- Hartvigsen et al. (2022) Thomas Hartvigsen, Saadia Gabriel, Hamid Palangi, Maarten Sap, Dipankar Ray, and Ece Kamar. 2022. ToxiGen: A Large-Scale Machine-Generated Dataset for Adversarial and Implicit Hate Speech Detection. In Proc. of ACL. ACL, Dublin, Ireland, 3309–3326.
- Husain and Uzuner (2021) Fatemah Husain and Ozlem Uzuner. 2021. A survey of offensive language detection for the Arabic language. ACM TALLIP 20, 1 (2021), 1–44.
- Husain and Uzuner (2022) Fatemah Husain and Ozlem Uzuner. 2022. Transfer Learning Across Arabic Dialects for Offensive Language Detection. In Proc. of IALP. 196–205.
- Ibrohim and Budi (2019a) Muhammad Okky Ibrohim and Indra Budi. 2019a. Translated vs non-translated method for multilingual hate speech identification in Twitter. International Journal on Advanced Science, Engineering and Information Technology 9, 4 (1 Jan. 2019), 1116–1123.
- Ibrohim and Budi (2019b) Muhammad Okky Ibrohim and Indra Budi. 2019b. Multi-label Hate Speech and Abusive Language Detection in Indonesian Twitter. In Proc. of the ALW Workshop. ACL, Florence, Italy, 46–57.
- Jahan and Oussalah (2023) Md Saroar Jahan and Mourad Oussalah. 2023. A systematic review of Hate Speech automatic detection using Natural Language Processing. Neurocomputing (2023), 126232.
- Jeong et al. (2022) Younghoon Jeong, Juhyun Oh, Jongwon Lee, Jaimeen Ahn, Jihyung Moon, Sungjoon Park, and Alice Oh. 2022. KOLD: Korean Offensive Language Dataset. In Proc. of EMNLP. ACL, Abu Dhabi, United Arab Emirates, 10818–10833.
- Jiang et al. (2022) Aiqi Jiang, Xiaohan Yang, Yang Liu, and Arkaitz Zubiaga. 2022. SWSR: A Chinese dataset and lexicon for online sexism detection. Online Social Networks and Media 27 (2022), 100182.
- Jiang and Zubiaga (2021) Aiqi Jiang and Arkaitz Zubiaga. 2021. Cross-Lingual Capsule Network for Hate Speech Detection in Social Media. In Proc. of ACM HT (Virtual Event, USA). ACM, New York, NY, USA, 217–223.
- Jiang and Zubiaga (2023) Aiqi Jiang and Arkaitz Zubiaga. 2023. SexWEs: Domain-Aware Word Embeddings via Cross-Lingual Semantic Specialisation for Chinese Sexism Detection in Social Media. Proc. of ICWSM 17, 1 (Jun. 2023), 447–458.
- Jin et al. (2020) Di Jin, Zhijing Jin, Joey Tianyi Zhou, and Peter Szolovits. 2020. Is bert really robust? a strong baseline for natural language attack on text classification and entailment. In Proc. of AAAI, Vol. 34. 8018–8025.
- Kapoor et al. (2019) Raghav Kapoor, Yaman Kumar, Kshitij Rajput, Rajiv Ratn Shah, Ponnurangam Kumaraguru, and Roger Zimmermann. 2019. Mind Your Language: Abuse and Offense Detection for Code-Switched Languages. Proc. of AAAI 33, 01 (Jul. 2019), 9951–9952.
- Kennedy et al. (2020) Brendan Kennedy, Xisen Jin, Aida Mostafazadeh Davani, Morteza Dehghani, and Xiang Ren. 2020. Contextualizing hate speech classifiers with post-hoc explanation. arXiv preprint arXiv:2005.02439 (2020).
- Khan et al. (2021) Muhammad Moin Khan, Khurram Shahzad, and Muhammad Kamran Malik. 2021. Hate Speech Detection in Roman Urdu. ACM Transactions on Asian and Low-Resource Language Information Processing 20, 1, Article 9 (Mar. 2021).
- Kiritchenko et al. (2021) Svetlana Kiritchenko, Isar Nejadgholi, and Kathleen C Fraser. 2021. Confronting abusive language online: A survey from the ethical and human rights perspective. Artificial Intelligence Research 71 (2021), 431–478.
- Koch et al. (2015) Gregory Koch, Richard Zemel, Ruslan Salakhutdinov, et al. 2015. Siamese neural networks for one-shot image recognition. In ICML deep learning workshop, Vol. 2. Lille.
- Kumar and Sachdeva (2022) Akshi Kumar and Nitin Sachdeva. 2022. Multi-input integrative learning using deep neural networks and transfer learning for cyberbullying detection in real-time code-mix data. Multimedia systems 28, 6 (2022), 2027–2041.
- Kumar et al. (2020) Ritesh Kumar, Atul Kr Ojha, Shervin Malmasi, and Marcos Zampieri. 2020. Evaluating aggression identification in social media. In Proc. of the TRAC Workshop. 1–5.
- Kumar et al. (2018) Ritesh Kumar, Aishwarya N. Reganti, Akshit Bhatia, and Tushar Maheshwari. 2018. Aggression-annotated Corpus of Hindi-English Code-mixed Data. In Proc. of LREC. ELRA, Miyazaki, Japan.
- Lai et al. (2023) Mirko Lai, Fabio Celli, Alan Ramponi, Sara Tonelli, Cristina Bosco, and Viviana Patti. 2023. Haspeede3 at evalita 2023: Overview of the political and religious hate speech detection task. In EVALITA.
- Lample et al. (2018a) Guillaume Lample, Alexis Conneau, Ludovic Denoyer, and Marc’Aurelio Ranzato. 2018a. Unsupervised machine translation using monolingual corpora only. In Proc. of ICLR.
- Lample et al. (2018b) Guillaume Lample, Alexis Conneau, Marc’Aurelio Ranzato, Ludovic Denoyer, and Hervé Jégou. 2018b. Word translation without parallel data. In Proc. of ICLR.
- Lee et al. (2021) Chanhee Lee, Kisu Yang, Taesun Whang, Chanjun Park, Andrew Matteson, and Heuiseok Lim. 2021. Exploring the Data Efficiency of Cross-Lingual Post-Training in Pretrained Language Models. Applied Sciences 11, 5 (2021).
- Lee et al. (2013) Dong-Hyun Lee et al. 2013. Pseudo-label: The simple and efficient semi-supervised learning method for deep neural networks. In ICML Workshop on challenges in representation learning, Vol. 3. Atlanta, 896.
- Liu et al. (2020) Yinhan Liu, Jiatao Gu, Naman Goyal, Xian Li, Sergey Edunov, Marjan Ghazvininejad, Mike Lewis, and Luke Zettlemoyer. 2020. Multilingual Denoising Pre-training for Neural Machine Translation. Transactions of the ACL 8 (2020), 726–742.
- Ljubešić et al. (2019) Nikola Ljubešić, Darja Fišer, and Tomaž Erjavec. 2019. The FRENK datasets of socially unacceptable discourse in Slovene and English. In Proc. of TSD. Springer, 103–114.
- Lundberg and Lee (2017) Scott M Lundberg and Su-In Lee. 2017. A Unified Approach to Interpreting Model Predictions. In Proc. of NeurIPS 30, I. Guyon, U. V. Luxburg, S. Bengio, H. Wallach, R. Fergus, S. Vishwanathan, and R. Garnett (Eds.). Curran Associates, Inc., 4765–4774.
- Mandl et al. (2021) Thomas Mandl, Sandip Modha, Anand Kumar M, and Bharathi Raja Chakravarthi. 2021. Overview of the HASOC Track at FIRE 2020: Hate Speech and Offensive Language Identification in Tamil, Malayalam, Hindi, English and German. In Proc. of FIRE (Hyderabad, India). ACM, New York, NY, USA, 29–32.
- Mandl et al. (2019) Thomas Mandl, Sandip Modha, Prasenjit Majumder, Daksh Patel, Mohana Dave, Chintak Mandlia, and Aditya Patel. 2019. Overview of the HASOC Track at FIRE 2019: Hate Speech and Offensive Content Identification in Indo-European Languages. In Proc. of FIRE (Kolkata, India). ACM, New York, NY, USA, 14–17.
- Markov et al. (2021) Ilia Markov, Nikola Ljubešić, Darja Fišer, and Walter Daelemans. 2021. Exploring Stylometric and Emotion-Based Features for Multilingual Cross-Domain Hate Speech Detection. In Proc. of the WASSA Workshop. ACL, Online, 149–159.
- Mathew et al. (2021) Binny Mathew, Punyajoy Saha, Seid Muhie Yimam, Chris Biemann, Pawan Goyal, and Animesh Mukherjee. 2021. Hatexplain: A benchmark dataset for explainable hate speech detection. In Proc. of AAAI, Vol. 35. 14867–14875.
- Mathur et al. (2018a) Puneet Mathur, Ramit Sawhney, Meghna Ayyar, and Rajiv Shah. 2018a. Did you offend me? Classification of Offensive Tweets in Hinglish Language. In Proc. of the ALW Workshop. ACL, Brussels, Belgium, 138–148.
- Mathur et al. (2018b) Puneet Mathur, Rajiv Shah, Ramit Sawhney, and Debanjan Mahata. 2018b. Detecting Offensive Tweets in Hindi-English Code-Switched Language. In Proc. of the SocialNLP Workshop. ACL, Melbourne, Australia, 18–26.
- Mishra et al. (2019) Pushkar Mishra, Helen Yannakoudakis, and Ekaterina Shutova. 2019. Tackling online abuse: A survey of automated abuse detection methods. arXiv preprint arXiv:1908.06024 (2019).
- Modha et al. (2021) Sandip Modha, Thomas Mandl, Gautam Kishore Shahi, Hiren Madhu, Shrey Satapara, Tharindu Ranasinghe, and Marcos Zampieri. 2021. Overview of the hasoc subtrack at fire 2021: Hate speech and offensive content identification in english and indo-aryan languages and conversational hate speech. In Proce. of FIRE. 1–3.
- Montariol et al. (2022) Syrielle Montariol, Arij Riabi, and Djamé Seddah. 2022. Multilingual Auxiliary Tasks Training: Bridging the Gap between Languages for Zero-Shot Transfer of Hate Speech Detection Models. In Findings of AACL-IJCNLP. ACL, Online only, 347–363.
- Mozafari et al. (2022) Marzieh Mozafari, Reza Farahbakhsh, and Noel Crespi. 2022. Cross-Lingual Few-Shot Hate Speech and Offensive Language Detection Using Meta Learning. IEEE Access 10 (2022), 14880–14896.
- Mubarak et al. (2017) Hamdy Mubarak, Kareem Darwish, and Walid Magdy. 2017. Abusive Language Detection on Arabic Social Media. In Proc. of the ALW Workshop. ACL, Vancouver, BC, Canada, 52–56.
- Mubarak et al. (2020) Hamdy Mubarak, Kareem Darwish, Walid Magdy, Tamer Elsayed, and Hend Al-Khalifa. 2020. Overview of OSACT4 Arabic Offensive Language Detection Shared Task. In Proc. of the OSACT Workshop. ELRA, Marseille, France, 48–52.
- Mubarak et al. (2021) Hamdy Mubarak, Ammar Rashed, Kareem Darwish, Younes Samih, and Ahmed Abdelali. 2021. Arabic Offensive Language on Twitter: Analysis and Experiments. In Proc. of the WANLP Workshop. ACL, Kyiv, Ukraine (Virtual), 126–135.
- Mulki et al. (2019) Hala Mulki, Hatem Haddad, Chedi Bechikh Ali, and Halima Alshabani. 2019. L-hsab: A levantine twitter dataset for hate speech and abusive language. In Proc. of the ALW Workshop. 111–118.
- Muti and Barrón-Cedeño (2022) Arianna Muti and Alberto Barrón-Cedeño. 2022. A Checkpoint on Multilingual Misogyny Identification. In Proc. of ACL SRW. ACL, Dublin, Ireland, 454–460.
- Nozza (2021) Debora Nozza. 2021. Exposing the limits of Zero-shot Cross-lingual Hate Speech Detection. In Proc. of ACL-IJCNLP. ACL, Online, 907–914.
- OpenAI (2023) OpenAI. 2023. GPT-4 Technical Report. arXiv:2303.08774 [cs.CL]
- Ousidhoum et al. (2019) Nedjma Ousidhoum, Zizheng Lin, Hongming Zhang, Yangqiu Song, and Dit-Yan Yeung. 2019. Multilingual and Multi-Aspect Hate Speech Analysis. In Proc. of EMNLP-IJCNLP. ACL, Hong Kong, China, 4675–4684.
- Ouyang et al. (2021) Xuan Ouyang, Shuohuan Wang, Chao Pang, Yu Sun, Hao Tian, Hua Wu, and Haifeng Wang. 2021. ERNIE-M: Enhanced Multilingual Representation by Aligning Cross-lingual Semantics with Monolingual Corpora. In Proc. of EMNLP. ACL, Online and Punta Cana, Dominican Republic, 27–38.
- Page et al. (2021) Matthew J. Page, Joanne E. McKenzie, Patrick M. Bossuyt, Isabelle Boutron, Tammy C. Hoffmann, Cynthia D. Mulrow, Larissa Shamseer, Jennifer M. Tetzlaff, Elie A. Akl, Sue E. Brennan, Roger Chou, Julie Glanville, Jeremy M. Grimshaw, Asbjørn Hróbjartsson, Manoj M. Lalu, Tianjing Li, Elizabeth W. Loder, Evan Mayo-Wilson, Steve McDonald, Luke A. McGuinness, Lesley A. Stewart, James Thomas, Andrea C. Tricco, Vivian A. Welch, Penny Whiting, and David Moher. 2021. The PRISMA 2020 statement: An updated guideline for reporting systematic reviews. International Journal of Surgery 88 (2021), 105906.
- Pamungkas et al. (2020) Endang Wahyu Pamungkas, Valerio Basile, and Viviana Patti. 2020. Misogyny Detection in Twitter: a Multilingual and Cross-Domain Study. Information Processing & Management 57, 6 (2020), 102360.
- Pamungkas et al. (2021a) Endang Wahyu Pamungkas, Valerio Basile, and Viviana Patti. 2021a. A joint learning approach with knowledge injection for zero-shot cross-lingual hate speech detection. Information Processing & Management 58, 4 (2021), 102544.
- Pamungkas et al. (2021b) Endang Wahyu Pamungkas, Valerio Basile, and Viviana Patti. 2021b. Towards multidomain and multilingual abusive language detection: a survey. Personal and Ubiquitous Computing 27, 1 (2021), 17–43.
- Pamungkas and Patti (2019) Endang Wahyu Pamungkas and Viviana Patti. 2019. Cross-domain and Cross-lingual Abusive Language Detection: A Hybrid Approach with Deep Learning and a Multilingual Lexicon. In Proc. of ACL SRW. ACL, Florence, Italy, 363–370.
- Pan and Yang (2010) Sinno Jialin Pan and Qiang Yang. 2010. A Survey on Transfer Learning. IEEE Trans. on Knowledge and Data Engineering 22, 10 (2010), 1345–1359.
- Pant and Dadu (2020) Kartikey Pant and Tanvi Dadu. 2020. Towards Code-switched Classification Exploiting Constituent Language Resources. In Proc. of AACL-IJCNLP SRW. ACL, Suzhou, China, 37–43.
- Paul et al. (2023) Sayanta Paul, Sriparna Saha, and Jyoti Prakash Singh. 2023. COVID-19 and cyberbullying: deep ensemble model to identify cyberbullying from code-switched languages during the pandemic. Multimedia Tools and Applications 82, 6 (2023), 8773–8789.
- Pavlopoulos et al. (2021) John Pavlopoulos, Jeffrey Sorensen, Léo Laugier, and Ion Androutsopoulos. 2021. SemEval-2021 Task 5: Toxic Spans Detection. In Proc. of the SemEval Workshop. ACL, Online, 59–69.
- Pelicon et al. (2021a) Andraž Pelicon, Ravi Shekhar, Matej Martinc, Blaž Škrlj, Matthew Purver, and Senja Pollak. 2021a. Zero-shot Cross-lingual Content Filtering: Offensive Language and Hate Speech Detection. In Proc.of the EACL Hackashop. ACL, Online, 30–34.
- Pelicon et al. (2021b) Andraž Pelicon, Ravi Shekhar, Blaž Škrlj, Matthew Purver, and Senja Pollak. 2021b. Investigating cross-lingual training for offensive language detection. PeerJ Computer Science 7 (2021), e559.
- Pereira-Kohatsu et al. (2019) Juan Carlos Pereira-Kohatsu, Lara Quijano-Sánchez, Federico Liberatore, and Miguel Camacho-Collados. 2019. Detecting and Monitoring Hate Speech in Twitter. Sensors 19, 21 (2019).
- Peters et al. (2018) Matthew E. Peters, Mark Neumann, Mohit Iyyer, Matt Gardner, Christopher Clark, Kenton Lee, and Luke Zettlemoyer. 2018. Deep Contextualized Word Representations. In Proc. of NAACL-HLT. ACL, New Orleans, Louisiana, 2227–2237.
- Phung and Cloos (2022) Tung Minh Phung and Jan Cloos. 2022. An exploratory experiment on Hindi, Bengali hate-speech detection and transfer learning using neural networks. arXiv preprint arXiv:2201.01997 (2022).
- Pikuliak et al. (2021) Matúš Pikuliak, Marián Šimko, and Mária Bieliková. 2021. Cross-lingual learning for text processing: A survey. Expert Systems with Applications 165 (2021), 113765.
- Pitenis et al. (2020) Zesis Pitenis, Marcos Zampieri, and Tharindu Ranasinghe. 2020. Offensive Language Identification in Greek. In Proc. of LREC. ELRA, Marseille, France, 5113–5119.
- Plaza et al. (2023) Laura Plaza, Jorge Carrillo-de Albornoz, Roser Morante, Enrique Amigó, Julio Gonzalo, Damiano Spina, and Paolo Rosso. 2023. Overview of EXIST 2023: sEXism Identification in Social NeTworks. In Proc. of ECIR. Springer, 593–599.
- Plaza-del arco et al. (2023) Flor Miriam Plaza-del arco, Debora Nozza, and Dirk Hovy. 2023. Respectful or Toxic? Using Zero-Shot Learning with Language Models to Detect Hate Speech. In Proc. of the WOAH workshop. ACL, Toronto, Canada, 60–68.
- Poletto et al. (2021) Fabio Poletto, Valerio Basile, Manuela Sanguinetti, Cristina Bosco, and Viviana Patti. 2021. Resources and benchmark corpora for hate speech detection: a systematic review. Language Resources and Evaluation 55 (2021), 477–523.
- Ptaszyński et al. (2018) Michał Ptaszyński, Gniewosz Leliwa, Mateusz Piech, and Aleksander Smywiński-Pohl. 2018. Cyberbullying Detection–Technical Report 2/2018, Department of Computer Science AGH, University of Science and Technology. arXiv preprint arXiv:1808.00926 (2018).
- Ptaszynski et al. (2019) Michal Ptaszynski, Agata Pieciukiewicz, and Paweł Dybała. 2019. Results of the poleval 2019 shared task 6: First dataset and open shared task for automatic cyberbullying detection in polish twitter. (2019).
- Ranasinghe and Zampieri (2020) Tharindu Ranasinghe and Marcos Zampieri. 2020. Multilingual Offensive Language Identification with Cross-lingual Embeddings. In Proc. of EMNLP. ACL, Online, 5838–5844.
- Ranasinghe and Zampieri (2021a) Tharindu Ranasinghe and Marcos Zampieri. 2021a. An Evaluation of Multilingual Offensive Language Identification Methods for the Languages of India. Information 12, 8 (2021).
- Ranasinghe and Zampieri (2021b) Tharindu Ranasinghe and Marcos Zampieri. 2021b. MUDES: Multilingual Detection of Offensive Spans. In Proc. of NAACL-HLT: Demonstrations. ACL, Online, 144–152.
- Ranasinghe and Zampieri (2021c) Tharindu Ranasinghe and Marcos Zampieri. 2021c. Multilingual Offensive Language Identification for Low-Resource Languages. ACM Transactions on Asian and Low-Resource Language Information Processing 21, 1, Article 4 (Nov. 2021), 13 pages.
- Ribeiro et al. (2016) Marco Ribeiro, Sameer Singh, and Carlos Guestrin. 2016. “Why Should I Trust You?”: Explaining the Predictions of Any Classifier. In Proc. of NAACL: Demonstrations. ACL, San Diego, California, 97–101.
- Rivaling (2021) Bilingual Word Embeddings Rivaling. 2021. Llod-driven bilingual word embeddings rivaling cross-lingual transformers in quality of life concept detection from french online health communities. In Proc. of of SEMANTiCS, Vol. 53. IOS Press, 89.
- Rizwan et al. (2020) Hammad Rizwan, Muhammad Haroon Shakeel, and Asim Karim. 2020. Hate-speech and offensive language detection in roman Urdu. In Proc. of EMNLP. 2512–2522.
- Rodríguez et al. (2021) Sebastián E. Rodríguez, Héctor Allende-Cid, and Héctor Allende. 2021. Detecting Hate Speech in Cross-Lingual and Multi-lingual Settings Using Language Agnostic Representations. In Proc. of CIARP, João Manuel R. S. Tavares, João Paulo Papa, and Manuel González Hidalgo (Eds.). Springer, Cham, 77–87.
- Rodríguez-Sánchez et al. (2021) Francisco Rodríguez-Sánchez, Jorge Carrillo-de Albornoz, Laura Plaza, Julio Gonzalo, Paolo Rosso, Miriam Comet, and Trinidad Donoso. 2021. Overview of exist 2021: sexism identification in social networks. Procesamiento del Lenguaje Natural 67 (2021), 195–207.
- Rodríguez-Sánchez et al. (2022) Francisco Rodríguez-Sánchez, Jorge Carrillo-de Albornoz, Laura Plaza, Adrián Mendieta-Aragón, Guillermo Marco-Remón, Maryna Makeienko, María Plaza, Julio Gonzalo, Damiano Spina, and Paolo Rosso. 2022. Overview of exist 2022: sexism identification in social networks. Procesamiento del Lenguaje Natural 69 (2022), 229–240.
- Romim et al. (2021) Nauros Romim, Mosahed Ahmed, Hriteshwar Talukder, and Md Saiful Islam. 2021. Hate speech detection in the bengali language: A dataset and its baseline evaluation. In Proc. of IJCACI. Springer, 457–468.
- Ross et al. (2017) Björn Ross, Michael Rist, Guillermo Carbonell, Benjamin Cabrera, Nils Kurowsky, and Michael Wojatzki. 2017. Measuring the reliability of hate speech annotations: The case of the european refugee crisis. arXiv preprint arXiv:1701.08118 (2017).
- R”̈ottger et al. (2022a) Paul R”̈ottger, Debora Nozza, Federico Bianchi, and Dirk Hovy. 2022a. Data-Efficient Strategies for Expanding Hate Speech Detection into Under-Resourced Languages. In Proc. of EMNLP. ACL, Abu Dhabi, United Arab Emirates, 5674–5691.
- R”̈ottger et al. (2022b) Paul R”̈ottger, Haitham Seelawi, Debora Nozza, Zeerak Talat, and Bertie Vidgen. 2022b. Multilingual HateCheck: Functional Tests for Multilingual Hate Speech Detection Models. In Proc. of the WOAH Workshop. ACL, Seattle, Washington (Hybrid), 154–169.
- Roy et al. (2023) Sarthak Roy, Ashish Harshavardhan, Animesh Mukherjee, and Punyajoy Saha. 2023. Probing LLMs for hate speech detection: strengths and vulnerabilities. arXiv preprint arXiv:2310.12860 (2023).
- Ryzhova et al. (2022) Anastasia Ryzhova, Dmitry Devyatkin, Sergey Volkov, and Vladimir Budzko. 2022. Training Multilingual and Adversarial Attack-Robust Models for Hate Detection on Social Media. Procedia Computer Science 213 (2022), 196–202.
- Salaam et al. (2022) Cesa Salaam, Franck Dernoncourt, Trung Bui, Danda Rawat, and Seunghyun Yoon. 2022. Offensive Content Detection via Synthetic Code-Switched Text. In Proc. of COLING. ICCL, Gyeongju, Republic of Korea, 6617–6624.
- Sanguinetti et al. (2020) Manuela Sanguinetti, Gloria Comandini, Elisa Di Nuovo, Simona Frenda, Marco Stranisci, Cristina Bosco, Tommaso Caselli, Viviana Patti, and Irene Russo. 2020. Haspeede 2@ evalita2020: Overview of the evalita 2020 hate speech detection task. EVALITA (2020).
- Sanguinetti et al. (2018) Manuela Sanguinetti, Fabio Poletto, Cristina Bosco, Viviana Patti, and Marco Stranisci. 2018. An Italian Twitter Corpus of Hate Speech against Immigrants. In Proc. of LREC. ELRA, Miyazaki, Japan.
- Satapara et al. (2023) Shrey Satapara, Prasenjit Majumder, Thomas Mandl, Sandip Modha, Hiren Madhu, Tharindu Ranasinghe, Marcos Zampieri, Kai North, and Damith Premasiri. 2023. Overview of the HASOC Subtrack at FIRE 2022: Hate Speech and Offensive Content Identification in English and Indo-Aryan Languages. In Proc. of FIRE (Kolkata, India). ACM, New York, NY, USA, 4–7.
- Schmidt and Wiegand (2017) Anna Schmidt and Michael Wiegand. 2017. A survey on hate speech detection using natural language processing. In Proc. of the SocialNLP Workshop. 1–10.
- Shekhar et al. (2020) Ravi Shekhar, Marko Pranjic, Senja Pollak, Andraz Pelicon, and Matthew Purver. 2020. Automating News Comment Moderation with Limited Resources: Benchmarking in Croatian and Estonian. Journal for Language Technology and Computational Linguistics (2020).
- Shi et al. (2022) Xiayang Shi, Xinyi Liu, Chun Xu, Yuanyuan Huang, Fang Chen, and Shaolin Zhu. 2022. Cross-lingual offensive speech identification with transfer learning for low-resource languages. Computers and Electrical Engineering 101 (2022), 108005.
- Sigurbergsson and Derczynski (2020) Gudbjartur Ingi Sigurbergsson and Leon Derczynski. 2020. Offensive Language and Hate Speech Detection for Danish. In Proc. of LREC. ELRA, Marseille, France, 3498–3508.
- Smith et al. (2017) Samuel L. Smith, David H. P. Turban, Steven Hamblin, and Nils Y. Hammerla. 2017. Offline bilingual word vectors, orthogonal transformations and the inverted softmax. In Proc. of ICLR.
- Sohn and Lee (2019) Hajung Sohn and Hyunju Lee. 2019. MC-BERT4HATE: Hate Speech Detection using Multi-channel BERT for Different Languages and Translations. In 2019 International Conference on Data Mining Workshops (ICDMW). 551–559.
- Song et al. (2021) Guizhe Song, Degen Huang, and Zhifeng Xiao. 2021. A Study of Multilingual Toxic Text Detection Approaches under Imbalanced Sample Distribution. Information 12, 5 (2021).
- Stamou et al. (2022) Vivian Stamou, Iakovi Alexiou, Antigone Klimi, Eleftheria Molou, Alexandra Saivanidou, and Stella Markantonatou. 2022. Cleansing & expanding the HURTLEX(el) with a multidimensional categorization of offensive words. In Proc. of the WOAH Workshop. ACL, Seattle, Washington (Hybrid), 102–108.
- Stappen et al. (2020) Lukas Stappen, Fabian Brunn, and Bj”̈orn Schuller. 2020. Cross-lingual zero-and few-shot hate speech detection utilising frozen transformer language models and AXEL. arXiv preprint arXiv:2004.13850 (2020).
- Strus̈ et al. (2019) Julia Maria Strus̈, Melanie Siegel, Josef Ruppenhofer, Michael Wiegand, and Manfred Klenner. 2019. Overview of GermEval Task 2, 2019 shared task on the identification of offensive language. In Proc. of KONVENS. München [u.a.], 352 – 363.
- Swami et al. (2018) Sahil Swami, Ankush Khandelwal, Vinay Singh, Syed Sarfaraz Akhtar, and Manish Shrivastava. 2018. A corpus of english-hindi code-mixed tweets for sarcasm detection. arXiv preprint arXiv:1805.11869 (2018).
- Tiţa and Zubiaga (2021) Teodor Tiţa and Arkaitz Zubiaga. 2021. Cross-lingual Hate Speech Detection using Transformer Models. arXiv preprint arXiv:2111.00981 (2021).
- Touvron et al. (2023) Hugo Touvron, Thibaut Lavril, Gautier Izacard, Xavier Martinet, Marie-Anne Lachaux, Timothée Lacroix, Baptiste Rozière, Naman Goyal, Eric Hambro, Faisal Azhar, et al. 2023. Llama: Open and efficient foundation language models. arXiv preprint arXiv:2302.13971 (2023).
- UNSPAHS (2019) UNSPAHS. 2019. UN Strategy and Plan of Action on Hate Speech. https://www.un.org/en/genocideprevention/hate-speech-strategy.shtml. Accessed: 2023-10-24.
- Van Hee et al. (2015) Cynthia Van Hee, Els Lefever, Ben Verhoeven, Julie Mennes, Bart Desmet, Guy De Pauw, Walter Daelemans, and Véronique Hoste. 2015. Automatic detection and prevention of cyberbullying. In Proc. of HUSO. IARIA, 13–18.
- Vashistha and Zubiaga (2021) Neeraj Vashistha and Arkaitz Zubiaga. 2021. Online Multilingual Hate Speech Detection: Experimenting with Hindi and English Social Media. Information 12, 1 (2021).
- Vidgen and Derczynski (2020) Bertie Vidgen and Leon Derczynski. 2020. Directions in abusive language training data, a systematic review: Garbage in, garbage out. Plos One 15, 12 (2020), e0243300.
- Vidgen et al. (2020) Bertie Vidgen, Tristan Thrush, Zeerak Waseem, and Douwe Kiela. 2020. Learning from the worst: Dynamically generated datasets to improve online hate detection. arXiv preprint arXiv:2012.15761 (2020).
- Vitiugin et al. (2021) Fedor Vitiugin, Yasas Senarath, and Hemant Purohit. 2021. Efficient Detection of Multilingual Hate Speech by Using Interactive Attention Network with Minimal Human Feedback. In Proc. of ACM WebSci (Virtual Event, United Kingdom). ACM, New York, NY, USA, 130–138.
- Wang and Zheng (2015) Dong Wang and Thomas Fang Zheng. 2015. Transfer learning for speech and language processing. In Proc. of APSIPA. IEEE, 1225–1237.
- Wang et al. (2023) Han Wang, Ming Shan Hee, Md Rabiul Awal, Kenny Tsu Wei Choo, and Roy Ka-Wei Lee. 2023. Evaluating GPT-3 Generated Explanations for Hateful Content Moderation. In Proc. of IJCAI. 6255–6263.
- Waseem and Hovy (2016) Zeerak Waseem and Dirk Hovy. 2016. Hateful Symbols or Hateful People? Predictive Features for Hate Speech Detection on Twitter. In Proc. of NAACL SRW. ACL, San Diego, California, 88–93.
- Wei et al. (2021) Jason Wei, Maarten Bosma, Vincent Y Zhao, Kelvin Guu, Adams Wei Yu, Brian Lester, Nan Du, Andrew M Dai, and Quoc V Le. 2021. Finetuned language models are zero-shot learners. arXiv preprint arXiv:2109.01652 (2021).
- Wiegand et al. (2018) Michael Wiegand, Melanie Siegel, and Josef Ruppenhofer. 2018. Overview of the germeval 2018 shared task on the identification of offensive language. (2018).
- Wulczyn et al. (2017) Ellery Wulczyn, Nithum Thain, and Lucas Dixon. 2017. Ex machina: Personal attacks seen at scale. In Proc. of WWW. 1391–1399.
- Xue et al. (2021) Linting Xue, Noah Constant, Adam Roberts, Mihir Kale, Rami Al-Rfou, Aditya Siddhant, Aditya Barua, and Colin Raffel. 2021. mT5: A Massively Multilingual Pre-trained Text-to-Text Transformer. In Proc. of NAACL-HLT. ACL, Online, 483–498.
- Yi and Zubiaga (2023) Peiling Yi and Arkaitz Zubiaga. 2023. Session-based cyberbullying detection in social media: A survey. Online Social Networks and Media 36 (2023), 100250.
- Yin and Zubiaga (2021) Wenjie Yin and Arkaitz Zubiaga. 2021. Towards generalisable hate speech detection: a review on obstacles and solutions. PeerJ Computer Science 7 (2021), e598.
- zahra El-Alami et al. (2022) Fatima zahra El-Alami, Said Ouatik El Alaoui, and Noureddine En Nahnahi. 2022. A multilingual offensive language detection method based on transfer learning from transformer fine-tuning model. Computer and Information Sciences 34, 8, Part B (2022), 6048–6056.
- Zampieri et al. (2019) Marcos Zampieri, Shervin Malmasi, Preslav Nakov, Sara Rosenthal, Noura Farra, and Ritesh Kumar. 2019. SemEval-2019 Task 6: Identifying and Categorizing Offensive Language in Social Media (OffensEval). In Proc. of the SemEval Workshop, Jonathan May, Ekaterina Shutova, Aurelie Herbelot, Xiaodan Zhu, Marianna Apidianaki, and Saif M. Mohammad (Eds.). ACL, Minneapolis, Minnesota, USA, 75–86.
- Zampieri et al. (2020) Marcos Zampieri, Preslav Nakov, Sara Rosenthal, Pepa Atanasova, Georgi Karadzhov, Hamdy Mubarak, Leon Derczynski, Zeses Pitenis, and Çağrı Çöltekin. 2020. SemEval-2020 Task 12: Multilingual Offensive Language Identification in Social Media (OffensEval 2020). In Proc. of the SemEval Workshop. ICCL, Barcelona (online), 1425–1447.
- Zhang et al. (2023) Tianhua Zhang, Hongyin Luo, Yung-Sung Chuang, Wei Fang, Luc Gaitskell, Thomas Hartvigsen, Xixin Wu, Danny Fox, Helen Meng, and James Glass. 2023. Interpretable unified language checking. arXiv preprint arXiv:2304.03728 (2023).
- Zhao et al. (2023) Wayne Xin Zhao, Kun Zhou, Junyi Li, Tianyi Tang, Xiaolei Wang, Yupeng Hou, Yingqian Min, Beichen Zhang, Junjie Zhang, Zican Dong, et al. 2023. A survey of large language models. arXiv preprint arXiv:2303.18223 (2023).
- Zhou et al. (2023) Li Zhou, Laura Cabello, Yong Cao, and Daniel Hershcovich. 2023. Cross-Cultural Transfer Learning for Chinese Offensive Language Detection. In Proc. of the C3NLP Workshop). ACL, Dubrovnik, Croatia, 8–15.
- Zia et al. (2022) Haris Bin Zia, Ignacio Castro, Arkaitz Zubiaga, and Gareth Tyson. 2022. Improving Zero-Shot Cross-Lingual Hate Speech Detection with Pseudo-Label Fine-Tuning of Transformer Language Models. Proc. of ICWSM 16, 1 (May 2022), 1435–1439.